### v1

# Annotating Ambiguous Images: General Annotation Strategy for Image Classification with Real-World Biomedical Validation on Vertebral Fracture Diagnosis

Abstract
While numerous methods exist to solve classification problems within curated datasets, these solutions often fall short in biomedical applications due to the biased or ambiguous nature of the data. These difficulties are particularly evident when inferring height reduction from vertebral data, a key component of the clinically-recognized Genant score. Although strategies such as semi-supervised learning, proposal usage, and class blending may provide some resolution, a clear and superior solution remains elusive. This paper introduces a flowchart of general strategy to address these issues. We demonstrate the application of this strategy by constructing a vertebral fracture dataset with over 300,000 annotations. This work facilitates the transition of the classification problem into clinically meaningful scores and enriches our understanding of vertebral height reduction.
Keywords:
Biomedical application annotation strategydata-centric Vertebral fracture
1 Introduction
The diagnosis of vertebral fractures is a critical field in the biomedical community. [15] states that the relative risk of death after a vertebral fracture is almost 9 times higher and that almost 20% of women will experience another fracture within 1 year after a vertebral fracture. The classification of vertebral fractures is primarily based on the height reduction of the vertebrae. Due to degenerative deformities, it is not easy to determine fractures on the height loss alone. This issue leads to inconsistent / ambiguous annotations of the data, which leads to inferior training of a neural network. In this paper, we present a strategy to produce high quality data for this task, independent of the network architecture used.
Our primary focus is on the application of vertebral fracture diagnosis. However, we will also present a general strategy for annotating and generating high-quality data for ambiguous classification tasks. Ultimately, we will circle back to the vertebral fracture diagnosis application to validate the effectiveness of this general approach.
In general, deep learning based methods yield state-of-the-art results in image classification [17]. However, their success is founded in large amounts of high quality data [3, 61]. While the amount of labeled data may be reduced with self- or semi-supervision for image classification, some high quality labeled data is still needed especially for domain-specific tasks [32]. However, a key challenge in annotating data is the reliability and consistency of annotators and thus their annotations. Annotations are typically created by humans and as such they may have different opinions or interpretations for difficult image classification tasks or when the images are ambiguous in nature. The literature concludes that a single annotation is not sufficient to the ensure high quality of labels in these cases [7, 13, 1, 50]. To address this challenge, soft labels as an average across multiple annotations have been proposed as a possible solution to this problem [45].
Despite the benefits of multiple annotations per image, annotation is costly and time-consuming. Thus, this approach is often not feasible for large datasets or even for smaller datasets where high-quality domain experts are required [25]. [44] proposes the use of proposal-guided annotations to overcome this issue. Proposal-guided annotation means that a preliminary trained network provides a rough estimate of the class for each image, which serves the annotators as guidance to the annotators in their decision making [8]. This guidance can help reduce the annotation time and increase the consistency and thus the overall quality [43, 46]
All of these challenges, low data quality and difficult labeling apply to vertebral fractures, but there are no well-established guidelines for precessing the data. Moreover, medical experts are in high-demand and cannot be burdened with annotating thousands of images multiple times. Our work makes several contributions to the field of data-centric deep learning for the application of vertebral fracture diagnosis and ambiguous image classification in general.
Our first core contributions are that we propose to unify the individual learning’s from the literature to a practical guide of how to annotate data, especially ambiguous data. Although previous work has explored the use of individual techniques,we provide a step-by-step strategy that covers all aspects of the annotation process. The goal of this guide is applicable to any image classification task to ensure high quality data.
Our second contribution is we verify the effectiveness of the proposed strategy by applying in on a dataset of vertebral fracture classification [35, 49]. The guidelines were defined before they were tested of the dataset, thus we can reflect and discuss the benefits of the proposed guidelines.
Importantly, we note that this is an application paper that aims to reflect and improve upon the existing literature, rather than introduce a novel method. By emphasizing the practical applications of our approach, we hope to facilitate the adoption of da ata-centric perspective in more research and applications, while contributing to the ongoing conversation about best practices in deep learning research.
2 Practical Guidelines: How to annotate ambiguous data?
Before outlining our practical approach, we will address the rationale for its necessity and explain the concept of ambiguous data to which these guidelines should be applied. Common labeling guidelines [40, 5] use hard-encoded labels (one correct class) and do not consider the issue that arises from the fact that people may disagree about the class for a given image. This can lead to the presence of ambiguous labels or, more generally, ambiguous data – an issue that is common in many real-world applications [52, 42, 23, 4, 22, 62]. The literature provides several reasons why one annotation per image is insufficient to capture the ambiguity of the data [54, 45, 44, 7, 13, 1, 50]. The hard label probability distribution for an image $x$ can be described as $P(L^{x}=\cdot)\in\{0,1\}^{K}$ with $K$ classes.
Davani et al. showed that these multiple annotations should then be averaged to generate a soft label as in [45, 19], as opposed to using a majority vote, which is not optimal for ambiguous cases [7]. In addition, label smoothing, which involves smoothing hard-coded labels with a constant factor, is an established method for improving the performance of networks [55, 26, 36]. A soft-label probability distribution for $x$ can be described as $P(L^{x}=\cdot)\in[0,1]^{K}$.
It is important to note that ambiguous data is not just an error [39], but a data property. The data is inherently uncertain or ambiguous and thus must be treated as such. Thus, in general, the soft ground truth distribution $P(L^{x}=\cdot)$ is unknown and can only be approximated for the verification of the predicted results. This research is in line with a growing movement of data-centric deep learning [34, 21, 60, 13], where the focus is on data rather than model building.
We give an overview about the practical guidelines in a flowchart in Figure 1. It is divided into five main steps: Defining the specifics of what, who, and how the annotation process will be performed, the process itself, and its post-processing.
Figure 1: Flowchart with guidelines how to annotate ambiguous data, best viewed in color.
2.1 Definition - What?
The first step is to precisely define the classification task. In particular, the $K$ classes must be defined for the given problem and data. Additionally, raw unlabeled data must be collected, downloaded or generated based on the use case. The complete unlabeled image data is called $X_{r}$.
The next step is to determine the number of images to label. We looked at use cases with several thousand to several tens of thousands of images because we believe this is a common use case in real-world applications such as  [4, 43, 56, 10, 41]. When hundreds of thousands or even millions of images are available, self-supervised techniques may be applicable [6, 12]. However, even in these cases, performance evaluation requires a smaller labeled dataset that can be labeled with our strategy. We define $X_{u}\subseteq X_{r}$ as the subset containing only the samples that should be labeled.
The research we focus on builds on common supervised or semi-supervised deep learning [51, 53, 29]. Thus, we need enough samples per class to be available after annotation. The formulation ”enough samples” depends on the use cases, but our recommendation based on previously reported results would be that each class represents at least 1% of the complete data and that there are several dozen images per class, so that enough samples are available for evaluation [45]. If too few images are available, additional considerations about few or zero-shot learning approaches have to be taken into account [58, 57, 33]. However, this is beyond the scope of this paper.
For each image $x$ in a small labeled dataset $x\in X_{l}\subseteq X_{u}$ at least one hard-encoded ground truth label is needed. This data is used for annotator evaluation (see subsection 2.2) as well as potential training data of a proposal generation network (see subsection 2.3). An initial size of about 20% of the original dataset size was beneficial in [44], but it can be smaller or larger depending on the use-case.
2.2 Definition - Who?
A central question for the annotation process is who will annotate the image. Various concepts have been suggested in the literature, including the use of a single user, the collaboration of several workers, or even crowd-sourcing [40]. Depending on the background of the annotators, they need to be trained or selected for the given task [31]. To evaluate the quality of annoations, two elements are required in addition to the small labeled dataset for evaluation $X_{l}$: a threshold of acceptable performance $\mu$ and a suitable annotation platform.
The threshold $\mu$ depends on the evaluation metrics used. Macro accuracy or F1-score are commonly used and can show a high correlation [28, 16, 2, 45, 24, 30]. The results in [46] indicate that a threshold $\mu$ of 60% to 80% is often reasonable enough if the annotations are aggregated in the end. When choosing the threshold, one should keep in mind that the quality of the labels in $X_{l}$ affects the maximum achievable scores. For example, due to intra- or inter-observer variability, a score higher than 80% may not be possible [47].
Based on the task definition in subsection 2.1, a few classes with many occurrences are expected, so Sagar et al. recommend using a browsing environment for the annotation process, where multiple images can be viewed side by side [40]. However, other annotation platforms such as [48] are possible. These platforms may introduce an additional overhead during the setup, but can dramatically increase the speed of annotation.
The annotators should be trained until they exceed the desired performance $\mu$ on the labeled data $X_{l}$ with the chosen platform.
2.3 Definition - How?
The key question in this section is whether or not to use proposals in the annotation process. This question is central due to the fact that proposals can reduce the cost of annotation by a multitude, but may introduce unwanted side effects [20, 44]. This side-effect / bias can be introduced by to the fact that annotators are more likely to accept a proposed class, resulting in a skewed label distribution $P(L^{x}=\cdot)$ towards the proposed class. For this decision process, several parameters need to be approximated or at least their way of approximating them later should be defined. These parameters are motivated by [44] and are the class confusion matrix $c$ and the expected speedup $S$. We will first describe how these variables are used and refer to the supplement for possible approximation techniques. The most important part is to understand, when specific variables are needed, how they could be approximated at this step or how they will be approximated if they are only needed later.
The decision of whether to use proposals is guided by two main questions from [44]. The first question concerns the acceptability of introducing a bias. Although introducing a bias may seem undesirable, it is a bias that has been approved by human annotators. In recommendation systems [8, 9, 14], this bias is often inherently accepted, and even when humans reach a consensus [35], they may influence each other. If the conditions allow for an acceptable bias, proposals should be utilized. If not, the expected speedup must be evaluated.
The benefit of proposals is greater with a higher speedup ($S$), as a higher speedup enables easier reversal of the introduced bias with more annotations. However, there is a trade-off point where the speedup is too small to justify the introduced bias, and it becomes unlikely that the bias can be recovered with an equal or smaller budget. In [44], this tradeoff point is defined to be around 3. If the speedup falls below this threshold, the recommendation is to avoid using proposals and resort to normal annotations. Conversely, if the speedup exceeds the threshold, proposals can be employed.
The class confusion matrix $c$ is used in post-processing (see subsection 2.5) to improve the final distribution and is not needed until then.
2.4 Annotation Process
This step involves the most work because it concentrates on creating the annotations. Several preparatory steps need to be taken, depending on whether the utilization of proposals is intended or not. Previous literature has shown that proposals improve the consistency of annotations [38, 8], especially the use of overclustering seems to be beneficial with ambiguous data [43]. Therefore, we recommend using DC3 proposals [46], which have been shown to perform better on real-world ambiguous data than common network predictions and work with many semi-supervised methods [51, 30, 53]. The initial labeled data $X_{l}$ (defined in subsection 2.1) can be used as training data for such an arbitrary neural network. The network should be trained until it reaches the same quality threshold $\mu$ as the annotators. Before starting the annotation process, two variables must be defined based on the decision in subsection 2.3: the number of annotations until consensus ($A_{c}$) and the total number of annotations ($A^{c}$). These variables play a crucial role in determining the trade-off point between achieving the best approximation of the underlying distribution $P(L^{x}=\cdot)$ and minimizing the annotation effort and associated costs.
$A^{c}$ represents the general number of annotations per image, reflecting the quality of the approximation. While some data in $X_{u}$ may exhibit ambiguity and result in disagreement among annotators, there is a portion of data where all annotators reach a consensus. Although the specific proportion of complete consensus ($p_{c}$) has not been reported, [46] reports the proportion of at least 95% consistent annotations ($\hat{p_{c}}$) ranging from about 50% to 80% across four different datasets. To reduce the total annotation costs, we propose annotating only those images that do not achieve complete consensus after $A_{c}$ annotations. This approach aims to identify images where complete consensus can be achieved with fewer annotations than $A^{c}$.
The values chosen for $A_{c}$ and $A^{c}$ depend on the specific classification problem and previous design choices. In the supplementary section, we discuss plausible numbers of annotations per hour ($\frac{a}{h}$) and potential 95% confidence guarantees for different combinations of $A_{c}$ and $A^{c}$.
For the case where $A_{c}=10$ and $A^{c}=50$, two main conclusions can be drawn. First, based on the previously reported $\frac{a}{h}$ in [46], it is likely that all annotations can be completed within 10 days by a single annotator. This workload seems manageable, even for smaller projects, especially when shared among multiple annotators. Second, considering the approximation of $P(L^{x}=\cdot)$ and its unknown true value, if we want to achieve an approximation within a margin of 0.15 around the unknown true value with at least 95% confidence, using $A_{c}=10$ and $A^{c}=50$ is a reasonable choice.
2.5 Post-Processing
As discussed in subsection 2.3, the use of proposals during the annotation process introduces a bias due to the default effect [20]. It is important to note that while this speeds up the annotation process and produces more consistent annotations, it may also introduce a bias that may not be acceptable. Schmarje et al. propose in [44] with CleverLabel a combination of classblending and bias correction to improve the approximated distribution, which should be applied if proposals were used during the annotation process. If no proposals were used to create the annotations, it is still recommended to improve the distribution $P(L^{x}=\cdot)$ by blending with the class distribution $c$.
3 Evaluation
The goal of this paper is to solve real-world application annotation tasks such as the medical task of vertrebral fracture classification. For this purpose, we will describe how the previously defined strategy can be applied to this task on the publicly available datasets Verse2019[35] and Verse2020[49]. An analysis of the benefits or shortfalls of the decisions is given in subsection 3.2.
3.1 Applying the Strategy
We use the Verse datasets for better reproducibility of our osteoporotic vertrebral fracture classification. As in [35], we consider only at thoracolumbar vertebrae because osteoporotic fractures occur mainly in this region of the spine, as shown in Figure 2. For the classification of the fractures, we use an adaptation of the most commonly used score, the semi-quantitative Genant score [11]. As shown in Figure 2, the score defines four classes, which are mainly guided by the height reduction of the vertebrae compared to its neighbors or unfractured vertebrae. However, degenerative deformities (e.g. short vertebral height) of up to 20% are ignored for the osteoporotic fracture classification. Partially, these degenerative cases can not be assessed in the dataset because the resolution is lower than in normal radiographs. As shown in Table 1, this results in 3,761 individual vertebrae with a high class imbalance towards class zero. Overall, we have at least 1.1% samples per class and can apply common semi- or supervised techniques.
Figure 2: Illustration of a spine and definition of height reduction classes
[…]
Figure 3: Illustration of different image modalities
Unlike to the original author of the Verse2019 [35] dataset, we will not use medical experts for the annotation and will use 2D projections of the original 3D CT data.
Our decision to use non-medical experts has two main reasons. First, we want to show that non-experts can successfully annotate data with the help of proposals. Second, we plan to generate tens of thousands of annotations, which would not be feasible with medical experts, who are in high demand. While we will not use medical experts for the majority of the annotations, we have worked closely with the radiologist at XXXX (censored for review). We discussed our design choices with them, particularly regarding the definition of the task and the representation of the vertebrae used. As mentioned above, we trained the hired workers on this specific task and ensured that they achieved a macro accuracy and F1 score of at least $\mu=0.6$ on the complete dataset with and without proposals for at least two scores before they could annotate the final scores.
We do not use the original 3D CT data (see Figure 3) for the annotation but rather a 2D projection of the central slices for each vertebrae (see Figure 3) and its segmentation (see Figure 3). Additionally, we show the vertebrae above and below the target vertebrae to allow an evaluation of the height reduction. A blended example is shown in Figure 3. There are three major benefits/reasons of using 2D projections over 3D CT data. First, the Genant scores were defined on 2D X-Ray and 2D projections of CT data are visually very similar to X-Ray and thus are better comparable. Second, we want to use the reported methods from the literature for the proposal generation [46], which were only tested on 2D data. While it should be theoretically possible to extend the used networks to 3D, such major structural changes could lead to an inferior result which would dilute the results, such experiments may be future work. Lastly, when it comes to diagnostics, radiologists mainly look at one central slice which we include and the major 3D structures can be detected on this projection.
We hired 5 workers to annotate vertebrae in a web-based environment developed for this evaluation. Public annotation platforms do not support the combination of clusters and classes that we want to evaluate in this paper. Details about the platform are given in the supplementary. As $X_{l}$ we used the partially available labels from [35].
In a medical consensus process in an expert panel, the proposals of different people are unified, but influences between people due to social dynamics cannot be excluded. Therefore, we argue that a bias that may be introduced due to human approved annotations is comparable to a consensus process and therefore acceptable. This means that we will use proposals during the annotation and $c$ and $S$ only need to be approximated after the annotation process based on the defined strategy. To allow a validation of our decisions, we will also raise annotations without proposals and thus we can approximate the class distributions $c$ and speedup $S$ based on these annotations. Additional approaches to approximate $S$ as well as the approximation of a data-specific offset $\delta$ (see [44]) are discussed in the supplement.
Based on the training annotations, we estimated the $\frac{a}{h}$ to be about 2,000 and 3,000 without and with proposals, respectively. These results are similar values previously reported in [46]. For $A_{c}=10$, $A^{c}=50$ and estimated $p_{c}=0.5$, this would result in about 250,000 annotations with and without proposals. Furthermore, based on the estimated $\frac{a}{h}$ of about 2,500, this resulted in a workload of about 20 hours per annotator which was within our given budget. So we chose the number of annotations at the recommended values of ${A_{c}=10,~{}A^{c}=50}$. Additionally, we planned to annotate the 3D data to evaluate the quality difference between 2D and 3D. We limited these annotations to the test set of the original Verse2019 dataset [35].
3.2 Analysis
The annotators completed more than 100 iterations of annotation, half of them with proposals and half without. However, one annotator did not meet the acceptance threshold $\mu$ and was excluded along with the training iterations, leaving 80 valid iterations or a total of about 250,000 annotations. Additionally, 11 iterations were performed using 3D data from the test dataset for comparison. The average annotation rate was calculated to be 3,259.36 $\frac{a}{h}$, which means that it takes about one hour to complete the 3,761 annotations with the additional technical overhead and pauses. The comparison of the resulting dataset-specific variables with previously reported values [46] is given in Table 1. We see that the dataset offset $\delta$ and the rate of near-consensus rate $\hat{p_{c}}$ are similar to previously reported results. It is important to note that $p_{c}=0.27$, which is significantly smaller, resulting in more data with $A^{c}$ number of annotations. It should be evaluated in the future, if instead of total consensus a high number of agreeing votes is sufficient to reduce the cost of the annotation process. The largest class is greater than in the previous reports and the speedup $S$ is smaller in Table 1. This relationship is theoretically justified because a proposal is less important since the majority of images already belongs to a class and thus defines some kind of a proposal.
name $\delta$ $\hat{p_{c}}$ classes largest class [%] $S$
Verse (ours) 0.1143 0.6833 4 90.11 1.1636
CIFA10-H 0.0973 0.7766 10 10.16 2.4665
MiceBone 0.3636 0.4775 3 70.48 1.4471
Plankton 0.5784 0.7368 10 30.37 4.4319
Turkey 0.2164 0.6863 3 75.95 1.4877
Table 1: Comparison of dataset-specific variables with previous reported values [44]$\delta$ is the data specific offset when annotating with proposals described in [44]. $S$ is the speedup between annotating with and without proposals. $\hat{p_{c}}$ describes the percentage of data with at least 95% consistent annotations.
$F1$
2D - all 0.6380 $\pm$ 0.0323
2D - no proposal 0.6297 $\pm$ 0.0345
2D - with proposal 0.6462 $\pm$ 0.0302
3D - no proposal 0.6243 $\pm$ 0.0370
$F1$
Sampled from Human* 0.69
Proposal Network 0.57
Wei et. al [59] 0.59
Ours* 0.63
Table 2: Comparison human and network performance on Verse2019 [35] – * based on ten random sampled annotations
A comparison of the annotator and network performance on the Verse2019 [35] test set is given in Table 2. When comparing the macro F1 scores in 2D between annotations performed with and without proposals and those with proposal-based guidance achieved a significantly higher F1 score of about 2%. This means that proposal-based guidance led to better annotations by the annotators. The average macro F1 score for 3D annotations shows no significant difference compared to the 2D annotations, meaning that the modality difference was of no to little importance to the annotators.
The marginal class distribution on all vertebrae from the test set between the ground truth and the majority vote for 2D annotations was quite similar, with percentages of 80.33, 10.38, 6.01, and 3.28 versus 80.33, 10.93, 5.74, and 3.01, respectively. In addition, only 17% of the majority votes deviated from the ground truth. Comparing the average uncertainty based on $P(L^{x}=\cdot)$ for concurring and dissenting majority votes with the ground truth, we find a higher annotator uncertainty of 24.20% for dissenting scores than for concurring scores at 7.78%. These results indicate that the annotators produced annotations of similar quality to the original test data, with higher uncertainties for potentially difficult or incorrect majority votes.
The best performing network on the validation data for the proposal generation did not surpass the set target threshold $\mu=0.6$ with a macro F1 of 0.57. However, it surpassed $\mu$ on the validation and training data and it was decided that the performance was close enough. The proposal network was a semi-supervised Mean-Teacher model [53] with overclustering and default hyperparameters from [46]. The performance of the model was significantly better when trained only on the mask (see Figure 3) with only the desired vertebra visible instead of all other 2D modalities. We attribute this difference to the fact that the model could be more easily confused with vertebrae that are not currently of interest and to the more complicated representation of information e.g. in the form of blended images. In terms of the performance of the network trained on our new annotated data, but evaluated on the original test ground truth, we found that more annotations sampled from the annotated data generally led to better performance. This sampling from the human annotation could achieve random peaks up to about 0.7 F1-score. A standard ResNet50 model [18] with default hyperparameters (see supplementary for details) and cross-entropy as loss function could achieve a score of 0.63, which is higher than the more complex proposal network (trained on the original training data) and previously reported values by Wei et al. for non-specialized networks [59]. While they show that they can achieve higher scores with more problem-specific loss functions, with our data-centric approach, we can achieve an improvement without changing the model or training purely from the input data. Future potential improvement to also use special loss functions and better backend-models non-withstanding.
method number of annotations
proposal blending correction 1 3 5 10
DC3 balanced yes 0.4481 $\pm$ 0.1957 (0.00%) 0.2425 $\pm$ 0.0741 (0.00%) 0.1832 $\pm$ 0.0095 (0.00%) 0.1615 $\pm$ 0.0129 (0.00%)
DC3 balanced no 1.0309 $\pm$ 0.1750 (130.07%) 0.2704 $\pm$ 0.0760 (11.53%) 0.2105 $\pm$ 0.0256 (14.85%) 0.4546 $\pm$ 0.4343 (181.49%)
DC3 no yes 1.7382 $\pm$ 0.1553 (287.91%) 0.2499 $\pm$ 0.0374 (3.08%) 0.2081 $\pm$ 0.0171 (13.57%) 0.2798 $\pm$ 0.1325 (73.25%)
no only blends no 0.2311 $\pm$ 0.0224 (-48.42%) 0.2394 $\pm$ 0.0930 (-1.27%) 0.1904 $\pm$ 0.0209 (3.88%) 0.1644 $\pm$ 0.0062 (1.79%)
no balanced no 0.8565 $\pm$ 0.4466 (91.14%) 0.1966 $\pm$ 0.0245 (-18.90%) 0.1678 $\pm$ 0.0224 (-8.44%) 0.1568 $\pm$ 0.0165 (-2.92%)
no no no 0.5578 $\pm$ 0.1259 (24.49%) 0.2451 $\pm$ 0.0127 (1.10%) 0.5435 $\pm$ 0.4285 (196.61%) 0.1898 $\pm$ 0.0380 (17.52%)
Table 3: Resulting KL across different methods and number of annotations – The first method is the recommendation based on our defined guidelines. The next two methods use proposals and the last three methods do not use proposals which means their annotation time is also slower in comparison to the recommendation. Thus, one should not compare with the same number of annotations but higher number of annotations for DC3 proposals because the higher number can be achieved potentially in the same or less time. The results are given as KL $\pm$ STD ( Relative Change in % with regard to recommended method)
In Table 3, we compare the network predictions with the approximated soft ground truth distribution ($P(L^{x}=\cdot)$) using the Kullback-Leibler-divergence [27]. Comparing the use of DC3 proposals with and without blending or correction, we see that the results are better with both improvement methods proposed in [44]. Thus, we can confirm that the application of CleverLabel (the combination of both improvements) should be done in the post-processing step. We can also confirm that applying balanced blending to annotations without proposals improves the results. Using only the class blending based on the majority class seems to work particularly well for this task. However, as the number of annotations increases, the benefit decreases and from 5 annotations onward, it becomes worse than the recommended method of our defined strategy. In our description of how to apply the proposed annotation strategy (see subsection 3.1), we stated that the introduction of a bias is not a major concern. Nevertheless, we can validate that the strategy leads to the best results even if a bias would not be acceptable. Based on the speedup of about 1.2, the recommendation would have been to not annotate with proposals and to use only balanced blending, which is the best method for multiple annotations. Theoretically, if we had a high speedup, we could compare a higher number of annotations with proposals to fewer annotations without proposals. For example, a speedup at the decision boundary of 3, 5 annotations with proposal and 3 annotations without proposal, we have a lower annotation cost based on the calculations form [44] (5 $\cdot$ 0.33 vs. 3) and have a lower KL score. Thus, even in the case of higher speedups, our strategy suggests the correct choice.
4 Conclusion
In conclusion, this study presents a general strategy for annotating and processing ambiguous images with verification in vertebral fracture diagnosis. Overall, the human performance analysis demonstrated the effectiveness of proposal-guided annotation, resulting in improved F1 scores and reduced training time. A neural network trained on newly annotated data achieved a higher classification performance on the original test set and can potentially more than halve the KL score on our new baseline. Based on the results for all possible scenarios, we were able to verify the effectiveness of our strategy based on the literature for the given use case. By providing practical guidance and verifying its application on a real-world vertebral fracture dataset, this research contributes to the creation of high-quality datasets and advances the field of data-centric deep learning.
Supplementary Material
Approximation of workload
We look at the expected number of annotations per hour done by one annotator as $\frac{a}{h}$. Depending on the expected portion of complete consensus of the dataset $p_{c}$ and the total dataset size $N=|X_{u}|$, we can calculate the expected annotation time in hours with $A_{N}=N\cdot(p_{c}\cdot A_{c}+(1-p_{c})\cdot A^{c})/\frac{a}{h}$. In Figure 4 the required annotation time $A_{N}$ in days for $N=10,000,A_{c}=10,A^{c}=50$ with varying $p_{c}$ is given as well as values reported in [46]. The reported values for $\frac{a}{h}$ are between 1,000 and 4,500 without proposals or 3,000 and 10,000 with proposals. Thus, even in the worst-case the expected workdays with proposals are less than 10 days, or two weeks, which should be feasible in most cases especially when multiple annotators are available. These values are examples and have of course to be adopted to the specific use-case.
Figure 4: Annotation time in days based on expected consensus ratio $p_{c}$ and annotations per hour $\frac{a}{h}$, reported values from [46] and with 95% agreeing votes threshold $\hat{p_{c}}$
Approximation of confidences
Considering the binary case of the annotation task for class $k$ for each image $x$, i.e. whether class $k$ is annotated for $x$ or not, and having a sufficient number of annotations $A$, the 95%-confidence interval can be calculated using the Wald confidence interval [37]. The interval is given by
$\begin{split}P(L^{x}=k)&\pm Z_{95\%}\cdot\sqrt{\frac{p_{k}(1-p_{k})}{A}}\end{split}$ (1)
with 95%-confidence value, which is 1.96, and $p_{k}$ is the unknown true GT probability for class $k$. The confidence interval is the largest for $p=0.5$. The formula can be rearranged to calculate the required annotations $A$ if a 95% confidence interval around the approximation $P(L^{x}=k)$ of width $W$ and margin $W/2$ is to be guaranteed.
$\begin{split}\frac{W}{2}&=Z_{95\%}\cdot\sqrt{\frac{p_{k}(1-p_{k})}{A}}\\ \frac{W^{2}}{4}&=Z_{95\%}^{2}\cdot\frac{p_{k}(1-p_{k})}{A}\\ A&=\frac{4\cdot Z_{95\%}^{2}\cdot p_{k}(1-p_{k})}{W^{2}}\\ \end{split}$ (2)
Figure 5 visualizes the required annotations $A$ for various $p$ and $W$. These calculations can guarantee with 95% confidence that the confidence interval has a width of $W\approx 1$ for $A=3$, $W\approx 0.62$ for $A=10$ and $W\approx 0.28$ for $A=50$. However, a high number of $A$ is indirectly assumed because the Wald confidence interval approximates the original binomial distribution with a normal distribution. This approximation is especially inaccurate for small numbers of $A$ and $p$ near 0 or 1, as pointed out by Lyman et al. [37]. They suggest to approximate the 95% confidence intervals with $((0.25)^{\frac{1}{A}},1)$ when $p$ is close to one. This would give a lower bound of the interval of $0.63$ for $A=3$, $0.87$ for $A=10$ and $0.97$ for $A=50$. All these calculations assume an idealized version of the annotators, which must be taken into account in their interpretation. For this reason, it will not be investigated further how the confidence intervals can be approximated more precisely, but conclude that a consensus decision is uncertain for $A=3$ annotations, but reasonable for ten or more annotations. Similarly, approximations of $p$ with a confidence interval width less than 0.3 are only reasonable for about 50 annotations.
Figure 5: Number of images for confidence width $W$, logarithmic scale
Additional minor considerations for Annotation Strategy
  • All decisions should typically be supported by domain experts or the end users which will use the labeled data in the end.
  • Constraints like collaboration modes and support for proposals should be considered. For example, if crowd-sourcing should be used for annotation, this restricts the appropriate platforms. Due to the fact if proposals will be used is decided later in subsection 2.3, it is advisable to ensure that these options are supported to avoid unnecessary backtracking.
  • If the training of the annotators is not successful it might be necessary to step back and rethink previous decision to allow a successful training. For example, special training with experts could be held, more detailed description of the task could be given and general feedback of annotators could be considered.
  • Proposals $\rho_{x}$ with a low ground truth probability ($P(L^{x}=\rho_{x})=0$) should be avoided and [46] indicates only up to 10% of all proposals should be of such poor quality to allow successful training. The distribution $P(L^{x}=\cdot)$ might not be known at this point but if available this should be checked in addition to the threshold $\mu$.
  • The variables $c,\delta$ and $S$ may be approximated on various ways. One possible approach is defined in [44] by annotating about 100 images at least 10 times for the estimation of $c$, $\delta$ and $P(L^{x}=\cdot)$. The speedup $S$ can for example be measured on a small subset with and without proposals during the training of the annotators. However, if the variables are not required during the decision process described above, they could also be approximated on the real annotation data or ignored entirely. For example, $c$, can easily be approximated on the complete dataset if no proposals are used during the annotation process and an approximation of $S$ is not required if a small introduction of a bias is acceptable. The variables could also be approximated with an educated guess either based on previous knowledge or experiments.
  • The question is how $A_{c},A^{c}$ should be selected. We concentrate on two major aspects, what is a feasible workload for the annotators and what guarantees can be given. See for more details the approximations in Approximation of workload and Approximation of confidences.
  • The blending step may not be necessary if the number of annotations is high (e.g. above 50), since the distribution is already approximated quite accurately.
Approximation of $\delta$
We used annotations created during the training of the annotators to estimate $\delta$. We randomly picked 20 examples without consensus from the annotations without proposals. Based on these samples, the proposal per image $p_{x}$, the probabilities $P(L_{x}=\rho_{x})$ with and without proposal and the formula for simulated proposal acceptance of [44], we calculated $\delta\approx 0.11$.
Used hyperparameters
The used hyperparameters are defined in Table 4. For more detail see the respective modified source code.
Parameter Proposal network Evaluation network
Architecture DC3[46] + Mean-Teacher[53] ResNet50[18]
Learning rate 0.03 0.1
Epochs 600$\dagger$ 60
Optimizer SGD SGD
Batch size 64 128
Weight Decay 0.0005 0.0005
Pretrained Weights N/A ImageNet
Source code [46] [45]
Table 4: Used hyperparameters for the proposal and evaluation network, $\dagger$ approximated based on number of image iterations, epochs not counted
Additional results of the learning effect
Regarding the learning effect of the annotators, the training iterations showed an increase in F1 scores over time. The improvement in F1 scores across all annotators between the training and the last three iterations was 6.79 without proposals and 9.19 with proposals. When comparing the benefit of DC3 proposals versus no proposals, the average improvement in F1 scores was 3.27 for the first three iterations and 0.41 for the last three iterations. The proposal still has in the end a benefit but the annotators can more successfully annotate event without them. Moreover, the introduction of proposal-based annotations resulted in a reduction in training time which also diminished over time. The first three iterations saw a reduction of 13.49 minutes, while the last three iterations experienced a decrease of 4.55 minutes. Overall, we can conclude that the proposals helped to improve the quality and reduce the annotation time. However, this effect was diminished overtime while the performance with and without proposals increase. Thus, we see a learning effect of the annotators.
Annotation platform
Figure 6 displays the used annotation website. It is important to note, that all images are displayed to the user in a grid. In the case without proposal (see Figure 6), the shown classes are mixed and all have to be annotated manually by the user. In the second case with proposal (see Figure 6), most of the images are of the same class and can be annotated simultaneously, while only the errors need to be corrected.
[…]
Figure 6: Screenshots of the annotation website on the MiceBone dataset from [45], The sidebar for data selection is hidden for data protection reasons and to keep the anonymity.

### v2

@@
-Annotating Ambiguous Images: General Annotation Strategy for Image Classification with Real-World Biomedical Validation on Vertebral Fracture Diagnosis
+Annotating Ambiguous Images: General Annotation Strategy for High-Quality Data with Real-World Biomedical Validation
 Abstract
-While numerous methods exist to solve classification problems within curated datasets, these solutions often fall short in biomedical applications due to the biased or ambiguous nature of the data.
-These difficulties are particularly evident when inferring height reduction from vertebral data, a key component of the clinically-recognized Genant score.
-Although strategies such as semi-supervised learning, proposal usage, and class blending may provide some resolution, a clear and superior solution remains elusive.
-This paper introduces a flowchart of general strategy to address these issues.
-We demonstrate the application of this strategy by constructing a vertebral fracture dataset with over 300,000 annotations.
-This work facilitates the transition of the classification problem into clinically meaningful scores and enriches our understanding of vertebral height reduction.
-Keywords:
-Biomedical application annotation strategydata-centric Vertebral fracture
+In the field of image classification, existing methods often struggle with biased or ambiguous data, a prevalent issue in real-world scenarios.
+Current strategies, including semi-supervised learning and class blending, offer partial solutions but lack a definitive resolution.
+Addressing this gap, our paper introduces a novel strategy for generating high-quality labels in challenging datasets.
+Central to our approach is a clearly designed flowchart, based on a broad literature review, which enables the creation of reliable labels.
+We validate our methodology through a rigorous real-world test case in the biomedical field, specifically in deducing height reduction from vertebral imaging.
+Our empirical study, leveraging over 250,000 annotations, demonstrates the effectiveness of our strategies decisions compared to their alternatives.
+Keywords: Data quality management, ambiguous data, annotation data application, data-centric AI, vertebral fracture
 1 Introduction
-The diagnosis of vertebral fractures is a critical field in the biomedical community.
-[15] states that the relative risk of death after a vertebral fracture is almost 9 times higher and that almost 20% of women will experience another fracture within 1 year after a vertebral fracture.
-The classification of vertebral fractures is primarily based on the height reduction of the vertebrae.
-Due to degenerative deformities, it is not easy to determine fractures on the height loss alone.
-This issue leads to inconsistent / ambiguous annotations of the data, which leads to inferior training of a neural network.
-In this paper, we present a strategy to produce high quality data for this task, independent of the network architecture used.
-Our primary focus is on the application of vertebral fracture diagnosis.
-However, we will also present a general strategy for annotating and generating high-quality data for ambiguous classification tasks.
-Ultimately, we will circle back to the vertebral fracture diagnosis application to validate the effectiveness of this general approach.
-In general, deep learning based methods yield state-of-the-art results in image classification [17].
-However, their success is founded in large amounts of high quality data [3, 61].
-While the amount of labeled data may be reduced with self- or semi-supervision for image classification, some high quality labeled data is still needed especially for domain-specific tasks [32].
-However, a key challenge in annotating data is the reliability and consistency of annotators and thus their annotations.
-Annotations are typically created by humans and as such they may have different opinions or interpretations for difficult image classification tasks or when the images are ambiguous in nature.
-The literature concludes that a single annotation is not sufficient to the ensure high quality of labels in these cases [7, 13, 1, 50].
-To address this challenge, soft labels as an average across multiple annotations have been proposed as a possible solution to this problem [45].
-Despite the benefits of multiple annotations per image, annotation is costly and time-consuming.
-Thus, this approach is often not feasible for large datasets or even for smaller datasets where high-quality domain experts are required [25].
-[44] proposes the use of proposal-guided annotations to overcome this issue.
-Proposal-guided annotation means that a preliminary trained network provides a rough estimate of the class for each image, which serves the annotators as guidance to the annotators in their decision making [8].
-This guidance can help reduce the annotation time and increase the consistency and thus the overall quality [43, 46]
-All of these challenges, low data quality and difficult labeling apply to vertebral fractures, but there are no well-established guidelines for precessing the data.
-Moreover, medical experts are in high-demand and cannot be burdened with annotating thousands of images multiple times.
-Our work makes several contributions to the field of data-centric deep learning for the application of vertebral fracture diagnosis and ambiguous image classification in general.
-Our first core contributions are that we propose to unify the individual learning’s from the literature to a practical guide of how to annotate data, especially ambiguous data.
-Although previous work has explored the use of individual techniques,we provide a step-by-step strategy that covers all aspects of the annotation process.
-The goal of this guide is applicable to any image classification task to ensure high quality data.
-Our second contribution is we verify the effectiveness of the proposed strategy by applying in on a dataset of vertebral fracture classification [35, 49].
-The guidelines were defined before they were tested of the dataset, thus we can reflect and discuss the benefits of the proposed guidelines.
-Importantly, we note that this is an application paper that aims to reflect and improve upon the existing literature, rather than introduce a novel method.
-By emphasizing the practical applications of our approach, we hope to facilitate the adoption of da ata-centric perspective in more research and applications, while contributing to the ongoing conversation about best practices in deep learning research.
+Deep learning methods are at the leading methods for image classification, dependent on the availability of substantial high-quality data (Beyer et al., 2020; Yun et al., 2021).
+While techniques like self- or semi-supervision can reduce the need for labeled data, high-quality labels remain a necessity, especially in domain-specific tasks (Liu et al., 2023).
+A critical challenge lies in the reliability and consistency of human annotations, especially for ambiguous or complex classification tasks.
+The consensus in current research suggests that single annotations are inadequate for ensuring label quality in such cases (Davani et al., 2022; Grossmann et al., 2022; Basile et al., 2021; Sharmanska et al., 2016).
+To mitigate this, the adoption of soft labels, derived from averaging multiple annotations, has been proposed (Schmarje et al., 2022a).
+Soft labels can capture the inherent data uncertainty, which is different from the model uncertainty often assumed by uncertainty estimation methods.
+However, obtaining multiple annotations per image is resource-intensive and impractical for large datasets or when expertise is scarce (Krizhevsky et al., 2012).
+Schmarje et al.
+(2023) introduces proposal-guided annotations, where a pre-trained network provides preliminary class estimates to guide annotators, thereby enhancing annotation efficiency and quality (Desmond et al., 2021; Schmarje et al., 2021a, 2022b).
+An illustrative example of this process is shown in Figure 1.
+In this paper, we present a comprehensive strategy for generating high-quality data in ambiguous classification scenarios.
+High quality data means that we improve the label quality in comparison to one annotation only.
+We validate our approach in vertebral fracture diagnosis, a field with significant challenges due to data ambiguity, which is ideal for testing our strategy’s efficacy.
+Figure 1: Illustration of the concept of hard and soft labels and how they can be created from annotations – The recommended process has three steps.
+In the first step, an image $x$ is selected for annotation.
+The unknown ground-truth distribution ($P(\hat{L}^{x}=\cdot)$) could either be soft or hard as shown by the examples in the lower half.
+During the annotation, multiple annotations are created either with or without proposal.
+A proposal means that one class is recommended during the annotation process.
+In the example, class B is proposed and 32 annotations are generated.
+The average across these annotations could already be used as an approximation of the soft-label $P(L^{x}=\cdot)$, however it might be biased towards the proposal since it is more likely to accept a proposal Schmarje et al.
+(2023).
+Our post-processing step enhances the approximated distribution ($P(L^{x}=\cdot)$)from the second step by reducing this bias.
+In the provided example, the probability of class B is reduced since it was most likely overestimated due to the used proposal of class B.
+The diagnosis of vertebral fractures holds critical importance in medicine.
+As noted by Haczynski and Jakimiuk (2001), vertebral fractures significantly increase mortality risks and recurrence rates.
+Classifying these fractures, primarily based on vertebral height reduction, is challenging due to degenerative changes, leading to inconsistent annotations and impacting neural network training.
+This scenario underscores the necessity of our approach, especially when domain experts are scarce.
+Our primary contribution lies in synthesizing literature insights into a practical, step-by-step annotation guide for ambiguous real-world data.
+While individual annotation techniques have been explored previously, our unified strategy addresses the entire annotation process comprehensively.
+We aim to make this approach universally applicable to various image classification tasks, ensuring high-quality data.
+Furthermore, we will release software guidelines to assist in creating high-quality annotations following our approach.
+Our second key contribution is empirically validating our strategy on a vertebral fracture dataset (Löffler et al., 2020; Sekuboyina et al., 2021).
+By defining the guidelines prior to dataset testing, we provide an objective evaluation, underscoring the real-world applicability of our method.
+By focusing on practical applications, this work seeks to promote a data-centric perspective in research and application, contributing to the discourse on data quality best practices.
 2 Practical Guidelines: How to Annotate Ambiguous Data?
-Before outlining our practical approach, we will address the rationale for its necessity and explain the concept of ambiguous data to which these guidelines should be applied.
-Common labeling guidelines [40, 5] use hard-encoded labels (one correct class) and do not consider the issue that arises from the fact that people may disagree about the class for a given image.
-This can lead to the presence of ambiguous labels or, more generally, ambiguous data – an issue that is common in many real-world applications [52, 42, 23, 4, 22, 62].
-The literature provides several reasons why one annotation per image is insufficient to capture the ambiguity of the data [54, 45, 44, 7, 13, 1, 50].
-The hard label probability distribution for an image $x$ can be described as $P(L^{x}=\cdot)\in\{0,1\}^{K}$ with $K$ classes.
+Before detailing our approach, it’s essential to understand why these guidelines are crucial, especially in the context of ambiguous data.
+Typical labeling guidelines (Sager et al., 2021; Chang et al., 2017) rely on definitive, hard-encoded labels and often overlook the disagreement among annotators about a given image’s class.
+Such discrepancies lead to ambiguous labels or, in broader terms, ambiguous data – a prevalent issue in numerous real-world applications (Tarling et al., 2021; Sambyal et al., 2022; Karimi et al., 2020; Brünger et al., 2019; Jiang et al., 2021; Zhang et al., 2023).
+The probability distribution of a hard label for an image $x$ can be represented as $P(L^{x}=\cdot),P(\hat{L}^{x}=\cdot)\in\{0,1\}^{K}$, where $K$ is the number of classes and $L^{x},\hat{L}^{x}$ are random variables mapping an image $x$ to its class probability.
+We distinguish between $L^{x}$ and $\hat{L}^{x}$, which represent the estimated and ground-truth probability distributions for image $x$, respectively.
+The literature underlines that a single annotation per image is inadequate for capturing data ambiguity (Uma et al., 2021; Schmarje et al., 2022a, 2023; Davani et al., 2022; Grossmann et al., 2022; Basile et al., 2021; Sharmanska et al., 2016), necessitating multiple annotations for high-quality data.
 Davani et al.
-showed that these multiple annotations should then be averaged to generate a soft label as in [45, 19], as opposed to using a majority vote, which is not optimal for ambiguous cases [7].
-In addition, label smoothing, which involves smoothing hard-coded labels with a constant factor, is an established method for improving the performance of networks [55, 26, 36].
-A soft-label probability distribution for $x$ can be described as $P(L^{x}=\cdot)\in[0,1]^{K}$.
-It is important to note that ambiguous data is not just an error [39], but a data property.
-The data is inherently uncertain or ambiguous and thus must be treated as such.
-Thus, in general, the soft ground truth distribution $P(L^{x}=\cdot)$ is unknown and can only be approximated for the verification of the predicted results.
-This research is in line with a growing movement of data-centric deep learning [34, 21, 60, 13], where the focus is on data rather than model building.
-We give an overview about the practical guidelines in a flowchart in Figure 1.
-It is divided into five main steps: Defining the specifics of what, who, and how the annotation process will be performed, the process itself, and its post-processing.
-Figure 1: Flowchart with guidelines how to annotate ambiguous data, best viewed in color.
-2.1 Definition - What?
-The first step is to precisely define the classification task.
-In particular, the $K$ classes must be defined for the given problem and data.
-Additionally, raw unlabeled data must be collected, downloaded or generated based on the use case.
-The complete unlabeled image data is called $X_{r}$.
-The next step is to determine the number of images to label.
-We looked at use cases with several thousand to several tens of thousands of images because we believe this is a common use case in real-world applications such as  [4, 43, 56, 10, 41].
-When hundreds of thousands or even millions of images are available, self-supervised techniques may be applicable [6, 12].
-However, even in these cases, performance evaluation requires a smaller labeled dataset that can be labeled with our strategy.
-We define $X_{u}\subseteq X_{r}$ as the subset containing only the samples that should be labeled.
-The research we focus on builds on common supervised or semi-supervised deep learning [51, 53, 29].
-Thus, we need enough samples per class to be available after annotation.
-The formulation ”enough samples” depends on the use cases, but our recommendation based on previously reported results would be that each class represents at least 1% of the complete data and that there are several dozen images per class, so that enough samples are available for evaluation [45].
-If too few images are available, additional considerations about few or zero-shot learning approaches have to be taken into account [58, 57, 33].
-However, this is beyond the scope of this paper.
-For each image $x$ in a small labeled dataset $x\in X_{l}\subseteq X_{u}$ at least one hard-encoded ground truth label is needed.
-This data is used for annotator evaluation (see subsection 2.2) as well as potential training data of a proposal generation network (see subsection 2.3).
-An initial size of about 20% of the original dataset size was beneficial in [44], but it can be smaller or larger depending on the use-case.
-2.2 Definition - Who?
-A central question for the annotation process is who will annotate the image.
-Various concepts have been suggested in the literature, including the use of a single user, the collaboration of several workers, or even crowd-sourcing [40].
-Depending on the background of the annotators, they need to be trained or selected for the given task [31].
-To evaluate the quality of annoations, two elements are required in addition to the small labeled dataset for evaluation $X_{l}$: a threshold of acceptable performance $\mu$ and a suitable annotation platform.
-The threshold $\mu$ depends on the evaluation metrics used.
-Macro accuracy or F1-score are commonly used and can show a high correlation [28, 16, 2, 45, 24, 30].
-The results in [46] indicate that a threshold $\mu$ of 60% to 80% is often reasonable enough if the annotations are aggregated in the end.
-When choosing the threshold, one should keep in mind that the quality of the labels in $X_{l}$ affects the maximum achievable scores.
-For example, due to intra- or inter-observer variability, a score higher than 80% may not be possible [47].
-Based on the task definition in subsection 2.1, a few classes with many occurrences are expected, so Sagar et al.
-recommend using a browsing environment for the annotation process, where multiple images can be viewed side by side [40].
-However, other annotation platforms such as [48] are possible.
-These platforms may introduce an additional overhead during the setup, but can dramatically increase the speed of annotation.
-The annotators should be trained until they exceed the desired performance $\mu$ on the labeled data $X_{l}$ with the chosen platform.
-2.3 Definition - How?
-The key question in this section is whether or not to use proposals in the annotation process.
-This question is central due to the fact that proposals can reduce the cost of annotation by a multitude, but may introduce unwanted side effects [20, 44].
-This side-effect / bias can be introduced by to the fact that annotators are more likely to accept a proposed class, resulting in a skewed label distribution $P(L^{x}=\cdot)$ towards the proposed class.
-For this decision process, several parameters need to be approximated or at least their way of approximating them later should be defined.
-These parameters are motivated by [44] and are the class confusion matrix $c$ and the expected speedup $S$.
-We will first describe how these variables are used and refer to the supplement for possible approximation techniques.
-The most important part is to understand, when specific variables are needed, how they could be approximated at this step or how they will be approximated if they are only needed later.
-The decision of whether to use proposals is guided by two main questions from [44].
-The first question concerns the acceptability of introducing a bias.
-Although introducing a bias may seem undesirable, it is a bias that has been approved by human annotators.
-In recommendation systems [8, 9, 14], this bias is often inherently accepted, and even when humans reach a consensus [35], they may influence each other.
-If the conditions allow for an acceptable bias, proposals should be utilized.
-If not, the expected speedup must be evaluated.
-The benefit of proposals is greater with a higher speedup ($S$), as a higher speedup enables easier reversal of the introduced bias with more annotations.
-However, there is a trade-off point where the speedup is too small to justify the introduced bias, and it becomes unlikely that the bias can be recovered with an equal or smaller budget.
-In [44], this tradeoff point is defined to be around 3.
-If the speedup falls below this threshold, the recommendation is to avoid using proposals and resort to normal annotations.
-Conversely, if the speedup exceeds the threshold, proposals can be employed.
-The class confusion matrix $c$ is used in post-processing (see subsection 2.5) to improve the final distribution and is not needed until then.
-2.4 Annotation Process
-This step involves the most work because it concentrates on creating the annotations.
-Several preparatory steps need to be taken, depending on whether the utilization of proposals is intended or not.
-Previous literature has shown that proposals improve the consistency of annotations [38, 8], especially the use of overclustering seems to be beneficial with ambiguous data [43].
-Therefore, we recommend using DC3 proposals [46], which have been shown to perform better on real-world ambiguous data than common network predictions and work with many semi-supervised methods [51, 30, 53].
-The initial labeled data $X_{l}$ (defined in subsection 2.1) can be used as training data for such an arbitrary neural network.
-The network should be trained until it reaches the same quality threshold $\mu$ as the annotators.
-Before starting the annotation process, two variables must be defined based on the decision in subsection 2.3: the number of annotations until consensus ($A_{c}$) and the total number of annotations ($A^{c}$).
-These variables play a crucial role in determining the trade-off point between achieving the best approximation of the underlying distribution $P(L^{x}=\cdot)$ and minimizing the annotation effort and associated costs.
-$A^{c}$ represents the general number of annotations per image, reflecting the quality of the approximation.
-While some data in $X_{u}$ may exhibit ambiguity and result in disagreement among annotators, there is a portion of data where all annotators reach a consensus.
-Although the specific proportion of complete consensus ($p_{c}$) has not been reported, [46] reports the proportion of at least 95% consistent annotations ($\hat{p_{c}}$) ranging from about 50% to 80% across four different datasets.
-To reduce the total annotation costs, we propose annotating only those images that do not achieve complete consensus after $A_{c}$ annotations.
-This approach aims to identify images where complete consensus can be achieved with fewer annotations than $A^{c}$.
-The values chosen for $A_{c}$ and $A^{c}$ depend on the specific classification problem and previous design choices.
-In the supplementary section, we discuss plausible numbers of annotations per hour ($\frac{a}{h}$) and potential 95% confidence guarantees for different combinations of $A_{c}$ and $A^{c}$.
-For the case where $A_{c}=10$ and $A^{c}=50$, two main conclusions can be drawn.
-First, based on the previously reported $\frac{a}{h}$ in [46], it is likely that all annotations can be completed within 10 days by a single annotator.
-This workload seems manageable, even for smaller projects, especially when shared among multiple annotators.
-Second, considering the approximation of $P(L^{x}=\cdot)$ and its unknown true value, if we want to achieve an approximation within a margin of 0.15 around the unknown true value with at least 95% confidence, using $A_{c}=10$ and $A^{c}=50$ is a reasonable choice.
-2.5 Post-Processing
-As discussed in subsection 2.3, the use of proposals during the annotation process introduces a bias due to the default effect [20].
-It is important to note that while this speeds up the annotation process and produces more consistent annotations, it may also introduce a bias that may not be acceptable.
+(2022) demonstrated that averaging multiple annotations to form a soft label, as proposed in (Schmarje et al., 2022a; Hemming et al., 2018), is superior to using a majority vote in ambiguous cases.
+Moreover, label smoothing, a technique that tempers hard-coded labels with a constant factor, has been established as a method for enhancing network performance (Vaswani et al., 2017; Krothapalli and Abbott, 2020; Lukasik et al., 2020).
+A soft-label probability distribution for image $x$ can be expressed as $P(L^{x}=\cdot),P(\hat{L}^{x}=\cdot)\in[0,1]^{K}$.
+It is crucial to recognize that ambiguous data is not merely an error, as suggested in (Park and Chung, 2022), but a characteristic of the data itself.
+Such data inherently possess uncertainty and should be treated accordingly.
+As such, the soft ground truth distribution $P(\hat{L}^{x}=\cdot)$ is generally unknown and can only be approximated for validating predicted results.
+This approach aligns with the growing emphasis on data-centric deep learning (Liu et al., 2021b; Jarrahi et al., 2022; Whang et al., 2023; Grossmann et al., 2022), which prioritizes data quality over model architecture.
+We present a comprehensive overview of our practical guidelines through a flowchart in Figure 2.
+The guidelines encompass five main steps: defining the specifics of what, who, and how the annotation process will be executed, the annotation process itself, and its post-processing.
+The subsequent sections provide insights into the pipeline, with detailed discussions in Appendix A and Appendix B.
+Figure 2: Flowchart with guidelines on how to annotate ambiguous data, best viewed in color.
+  1. 1.
+    Definition - What?
+This stage outlines the classification task, specifying the classes ($K$) and gathering raw, unlabeled image data ($X_{r}$).
+A crucial element is selecting a representative subset ($X_{u}$) for annotation.
+The selection of $X_{u}$ emphasizes the importance of having an adequate sample size per class for effective training and evaluation or considers the use of self-supervised techniques for datasets without enough representative information.
+The definition of $X_{l}$ describes a smaller, precisely labeled dataset for evaluating annotators and training proposal generation networks.
+  2. 2.
+    Definition - Who?
+    This stage focuses on identifying suitable annotators for image labeling, ranging from individuals to crowdsourcing.
+It is important to train the annotators according to the task’s specific needs and setting a quality threshold for annotations ($\mu$), typically recommended between 60% and 80%.
+For classes with higher occurrence rates, a browsing environment for annotation is advised to improve process efficiency.
+  3. 3.
+    Definition - How?
+    In this stage, the decision to include proposals in the annotation process is discussed.
+Proposals can accelerate annotation but may introduce bias as discussed in (Schmarje et al., 2023).
+Key considerations include the acceptability of bias, the extent of speedup ($S$)), and a class confusion matrix ($c$) for later processing.
+Proposals are recommended if the bias is manageable and the speedup significant (typically above a threshold of 3), otherwise, standard annotations are preferred.
+  4. 4.
+    Annotation Process
+    This phase involves conducting the annotations, with or without proposals.
+For ambiguous data, overclustering (Schmarje et al., 2021b, a) is advised, and DC3 (Schmarje et al., 2022b) proposals are recommended for better performance.
+We propose to separate the annotations process into the number of annotations needed for early consensus ($A_{\text{cons}}$) and the total annotations for difficult cases($A^{\text{cons}}$).
+This separation balances ground-truth accuracy against the effort and cost of obtaining annotations.
+  5. 5.
+    Post-Processing
+    The concluding phase addresses the bias potentially introduced by proposals in the annotation process.
+CleverLabel (Schmarje et al., 2023), combining class blending and bias correction, is proposed for refining the approximated distribution if proposals were used.
+Even without proposals, blending the distribution with the class distribution ($c$) is recommended to enhance it.
+3 Evaluation
+Our objective is to address the challenges of annotating real-world application tasks, such as classifying vertebral fractures in medical images.
+To demonstrate the practicality of our proposed strategy, we apply it to the task using publicly available datasets Verse2019(Löffler et al., 2020) and Verse2020(Sekuboyina et al., 2021).
+3.1 Applying the Strategy
+The Verse datasets were chosen for their reproducibility and suitability in validating our workflow for classifying osteoporotic vertebral fractures.
+Following Sekuboyina et al.
+(2021); Löffler et al.
+(2020), we focus on thoracolumbar vertebrae, the primary site for osteoporotic fractures, as illustrated in Figure 4.
+Our classification approach is based on an adapted version of the Genant semi-quantitative score (Genant et al., 1996), which categorizes fractures into four classes based on the degree of height reduction in vertebrae compared to their neighbors.
+However, we exclude degenerative deformities, such as minor height reductions (up to 20%), which are often indistinguishable in the given CT images due to lower resolution compared to standard radiographs.
+As detailed in Figure 4, this results in a dataset of 3,761 individual vertebrae, with a notable class imbalance skewed towards class zero.
+Figure 3: Illustration of a spine and definition of height reduction classes
+Name $\delta$ $\hat{p_{c}}$ Classes Largest Class [%] $S$
+Verse (Ours) 0.1143 0.6833 4 90.11 1.1636
+CIFAR10-H 0.0973 0.7766 10 10.16 2.4665
+MiceBone 0.3636 0.4775 3 70.48 1.4471
+Plankton 0.5784 0.7368 10 30.37 4.4319
+Turkey 0.2164 0.6863 3 75.95 1.4877
+Figure 4: Comparison of dataset-specific variables with previously reported values (Schmarje et al., 2023)
+In contrast to the original authors of the Verse2019 dataset (Löffler et al., 2020), we employ non-medical experts for annotation and utilize 2D projections from the 3D CT data.
+The rationale and specifics of this approach are elaborated in the appendix Appendix C.
+Further details regarding data preprocessing, annotators, annotation platform, approximated variables, and used hyperparameters are provided in Appendix D.
+3.2 Analysis
+Our annotators completed over 100 iterations, half with proposals and half without.
+However, due to the failure of one annotator to meet the acceptance threshold $\mu$, their annotations, including training iterations, were excluded, leaving us with 80 valid iterations or about 250,000 annotations.
+An additional 11 iterations employed 3D data from the test dataset for comparative analysis.
+The average annotation rate was calculated at 3,259.36 annotations per hour, indicating that approximately two hours are needed to complete the 3,761 annotations, accounting for technical overhead and breaks.
+Comparison of dataset-specific variables with those in previous studies (Schmarje et al., 2022b) is provided in Figure 4.
+$\delta$ represents the data-specific offset when annotating with proposals as described in (Schmarje et al., 2023) and represents how strongly annotators are influenced in their decision by a proposal.
+0 means no influence while 1 means following the proposal completely.
+$S$ denotes the speedup between annotating with and without proposals.
+$\hat{p_{c}}$ indicates the percentage of data with at least 95% consistent annotations.
+Notably, the dataset offset $\delta$ and near-consensus rate $\hat{p_{c}}$ are similar to prior reports.
+We observed a larger dominant class and a smaller speedup $S$ in our dataset, as shown in Figure 4.
+This finding is theoretically consistent with the notion that a proposal is less impactful when the majority of images already belong to a specific class, as is the case with the high class imbalance towards class 0 in the Verse dataset.
+Our human annotator analysis indicated comparable quality to original test data, with increased uncertainties in cases of challenging or incorrect majority votes.
+Further details are provided in subsection E.2.
+Evaluation of Network Performance with Hard Labels
+Firstly, we evaluate the performance of our approach on the originally defined test data by (Löffler et al., 2020).
+Our best performing network for proposal generation achieved a macro F1 score of 0.57.
+This network was a semi-supervised Mean-Teacher model (Tarvainen and Valpola, 2017) with overclustering and default hyperparameters from (Schmarje et al., 2022b) and was trained on the original annotations on the subset $X_{l}$ as defined in Appendix C.
+A standard ResNet50 model (He et al., 2016) with default hyperparameters (details in the supplementary) and cross-entropy loss function attained a score of 0.63 if it was trained not on our newly annotated data.
+This outperformed the more complex proposal network and prior results by Wei et al.
+(2022) for non-specialized networks.
+Our data-centric approach enabled these improvements without model modifications, purely based on input data quality.
+Future research could explore further enhancements by incorporating specialized loss functions and advanced backend models into our workflow.
+Table 1: Results of KL divergence across different methods and number of annotations – The first method follows our recommended guidelines.
+The next two methods utilize proposals, while the last three do not, leading to a slower annotation time compared to the recommendation.
+Therefore, when comparing, one should consider a higher number of annotations for DC3 proposals, as they could potentially be achieved in the same or less time.
+The results are presented as KL $\pm$ STD (Relative Change in % compared to the recommended method).
+Method Number of Annotations
+Proposal Blending Correction 1 3 5 10
+DC3 Balanced Yes 0.4481 $\pm$ 0.1957 (0.00%) 0.2425 $\pm$ 0.0741 (0.00%) 0.1832 $\pm$ 0.0095 (0.00%) 0.1615 $\pm$ 0.0129 (0.00%)
+DC3 Balanced No 1.0309 $\pm$ 0.1750 (130.07%) 0.2704 $\pm$ 0.0760 (11.53%) 0.2105 $\pm$ 0.0256 (14.85%) 0.4546 $\pm$ 0.4343 (181.49%)
+DC3 No Yes 1.7382 $\pm$ 0.1553 (287.91%) 0.2499 $\pm$ 0.0374 (3.08%) 0.2081 $\pm$ 0.0171 (13.57%) 0.2798 $\pm$ 0.1325 (73.25%)
+No Only Blends No 0.2311 $\pm$ 0.0224 (-48.42%) 0.2394 $\pm$ 0.0930 (-1.27%) 0.1904 $\pm$ 0.0209 (3.88%) 0.1644 $\pm$ 0.0062 (1.79%)
+No Balanced No 0.8565 $\pm$ 0.4466 (91.14%) 0.1966 $\pm$ 0.0245 (-18.90%) 0.1678 $\pm$ 0.0224 (-8.44%) 0.1568 $\pm$ 0.0165 (-2.92%)
+No No No 0.5578 $\pm$ 0.1259 (24.49%) 0.2451 $\pm$ 0.0127 (1.10%) 0.5435 $\pm$ 0.4285 (196.61%) 0.1898 $\pm$ 0.0380 (17.52%)
+Evaluation of Network Performance with Soft Labels
+In Table 1, we compare the network predictions on the improved distributions with blending and correction (($P(L^{x}=\cdot)$) against the approximated soft ground truth distribution ($P(\hat{L}^{x}=\cdot)$) using Kullback-Leibler divergence (Kullback and Leibler, 1951).
+Comparing DC3 proposals with and without the proposed blending and correction steps in (Schmarje et al., 2023), we find improved results using both methods.
+Therefore, we confirm the efficacy of CleverLabel (a combination of these improvements) in the post-processing phase.
+Additionally, we can validate that applying balanced blending to annotations without proposals enhances results.
+Particularly, class blending based on the majority class proved effective for this task (Only Blends).
+However, the benefit diminishes with more than five annotations, at which point it becomes less effective than our strategy’s recommended method.
+In applying our proposed annotation strategy (subsection 3.1), we noted that bias introduction was not a major concern.
+Yet, our results affirm that the strategy leads to optimal outcomes even in scenarios where bias is unacceptable.
+Given a speedup of about 1.2, our flow chart would recommend to annotate without proposals, using only balanced blending, which emerged as the most effective method for multiple annotations.
+Theoretically, with higher speedups, we could compare a larger number of annotations with proposals to fewer annotations without proposals.
+For instance, at a decision boundary speedup of 3, five annotations with a proposal have a lower annotation cost than three without proposal (5 $\cdot$ 0.33 vs.
+3) and yield a reduced KL score.
+Hence, our strategy recommends the correct approach even in scenarios with higher speedups.
+3.3 Limitations and Future Work
+The use of a single dataset for validating our strategy may appear as a limitation, and the reliance on existing literature could be seen as a weakness.
+However, we contend that these aspects are, in fact, complementary.
+Our strategy, derived from a synthesis of recent research, leverages verified methodologies to ensure the validity of the overall approach, rather than just its individual components.
+Moreover, all potential decision paths in the graph were empirically tested.
+Thus, we believe the validation using the Verse datasets is adequate, chosen for their real-world relevance and abundance of ambiguous data, necessitating complex annotations and yielding insightful results.
+While the Verse datasets are smaller and less widely used than ImageNet (Russakovsky et al., 2015) or CIFAR10 (Krizhevsky and Hinton, 2009), their advantage lies in providing non-curated, class-distinct datasets.
+Although curated datasets like ImageNet include ambiguous classes (Beyer et al., 2020; Vasudevan et al., 2022; Peterson et al., 2019), they do not exhibit the level of noise and ambiguity found in real-world data as demonstrated in (Schmarje et al., 2021a, 2022a).
+The requirement for multiple annotations under various setups for this study necessitated over 250,000 annotations for roughly 4,000 images, making such an extensive evaluation unfeasible for larger datasets at present.
+While it is easier to annotate any classification task with our strategy, we still need many more annotations for evaluation, and thus evaluation is currently limited to smaller datasets.
+Future research aims to apply and validate this approach on a larger scale, and we invite fellow researchers to join us in this endeavor.
+4 Conclusion
+This study introduces a comprehensive strategy for annotating and processing ambiguous data, with a specific focus on vertebral fracture diagnosis.
+Our analysis of human annotator performance underscored the efficacy of proposal-guided annotation, which not only enhanced F1 scores but also streamlined the training process.
+A neural network trained on the newly annotated data exhibited superior classification performance on the original test set, potentially halving the Kullback-Leibler (KL) divergence score compared to our new baseline.
+Through thorough evaluation across all possible scenarios, we have substantiated the effectiveness of our strategy, rooted in literature, for this specific application.
+By providing practical guidelines and demonstrating their successful application in a real-world context, this research significantly contributes to the creation of high-quality datasets and the advancement of data-centric deep learning.
+5 Reproducibility Statement
+To ensure reproducibility, we have documented and made accessible all essential resources of this study.
+Our source code, data processing methods, model training, and evaluation procedures, along with a model card, are provided for transparency.
+We utilized the publicly available Verse’19 (Löffler et al., 2020) and Verse’20 (Sekuboyina et al., 2021) datasets, concentrating on thoracolumbar vertebrae for classifying osteoporotic fractures.
+Detailed preprocessing instructions are included to help with data reconstruction.
+The human annotation process was executed on a web-based platform, employing non-medical expert annotators who received specific training for this task.
+This project entailed over 250,000 annotations across approximately 4,000 images.
+To minimize variability inherent in human annotation, we established comprehensive guidelines and criteria for a consistent approach.
+For model training, we adopted a semi-supervised Mean-Teacher (Tarvainen and Valpola, 2017) model with overclustering, utilizing a ResNet50 architecture.
+Key hyperparameters included a learning rate of 0.03 for the proposal network and 0.1 for the evaluation network, batch sizes of 64 and 128, and a weight decay of 0.0005.
+The proposal network underwent training for approximately 600 epochs, while the evaluation network was trained for about 60 epochs.
+Acknowledgments
+We acknowledge funding of L.
+Schmarje by the ARTEMIS project (Grant number 01EC1908E) funded by the Federal Ministry of Education and Research (BMBF, Germany).
+We further acknowledge the funding of V.
+Grossmann by the Marispace-X project (grant number 68GX21002E), both funded by the Federal Ministry for Economic Affairs and Climate Action (BMWK, Germany).
+Appendix A Full Explanation of Flowchart
+A.1 Definition - What?
+The first crucial step involves defining the classification task.
+Specifically, this entails identifying the $K$ classes relevant to the given problem and dataset.
+Additionally, gathering or generating raw, unlabeled data, denoted as $X_{r}$, is necessary based on the use case.
+The next step involves deciding the number of images to label.
+We consider use cases with several thousand to tens of thousands of images, a common scenario in real-world applications (Brünger et al., 2019; Schmarje et al., 2021a; Volkmann et al., 2021; Durden et al., 2016; Saleh et al., 2020).
+For datasets with hundreds of thousands or millions of images, self-supervised techniques like SimCLRv2 or OmniMAE (Chen et al., 2020; Girdhar et al., 2023) may be applicable.
+However, even these cases require a smaller labeled dataset for performance evaluation, which can be labeled using our strategy.
+We denote $X_{u}\subseteq X_{r}$ as the subset containing samples to be labeled.
+Our focus is on common supervised or semi-supervised deep learning approaches (Sohn et al., 2020; Tarvainen and Valpola, 2017; Lee, 2013).
+Therefore, it’s imperative to ensure enough samples per class are available post-annotation.
+The term ”enough samples” varies depending on the use case.
+Based on previous results (Schmarje et al., 2022a), each class should represent at least 1% of the complete data, with several dozen images per class for robust evaluation.
+In cases of limited images, additional considerations for few or zero-shot learning approaches are necessary (Wang et al., 2020, 2019; Liu et al., 2021a), although this extends beyond our paper’s scope.
+For each image $x$ in a small labeled dataset $x\in X_{l}\subseteq X_{u}$, at least one hard-encoded ground truth label is necessary.
+This data is crucial for annotator evaluation (see subsection A.2) and potentially for training a proposal generation network (see subsection A.3).
+An initial dataset size of about 20% of the total dataset size was beneficial in (Schmarje et al., 2023), though the exact size may vary depending on the specific use-case.
+A.2 Definition - Who?
+The choice of annotators is a pivotal aspect of the annotation process.
+Various approaches include using a single user, collaborating with several workers, or employing crowd-sourcing methods (Sager et al., 2021).
+The annotators’ background dictates the extent of training or selection necessary for the task (Liao et al., 2021).
+For evaluating annotation quality, two elements are essential alongside the small labeled dataset $X_{l}$: a performance threshold $\mu$ and an appropriate annotation platform.
+The threshold $\mu$ should align with the evaluation metrics used, such as macro accuracy or F1-score, which often show high correlation (Laine and Aila, 2017; He et al., 2022; Berthelot et al., 2019; Schmarje et al., 2022a; Knausgård et al., 2021; Li et al., 2020).
+(Schmarje et al., 2022b) suggests that a threshold between 60% and 80% is typically adequate if annotations are aggregated.
+However, the label quality in $X_{l}$ influences the maximum achievable scores; for instance, scores higher than 80% might be unattainable due to intra- or inter-observer variability (Schmarje et al., 2019).
+Based on the task definition (see subsection A.1), a browsing environment is recommended for tasks with a few classes having many occurrences, allowing for side-by-side image comparison (Sager et al., 2021).
+Alternative platforms like (Schröder et al., 2020) are also viable.
+Although these platforms may require setup overhead, they can significantly expedite the annotation process.
+Annotators should be trained to exceed the performance threshold $\mu$ on the labeled data $X_{l}$ using the chosen platform.
+Another concern might be the annotator calibration, which is the alignment of annotations between the annotators.
+We actually do not require such alignment, because we measure these differences between the annotators.
+However, if an easy rule allows better alignment of annotators and yields less ambiguous results this direction can be pursued.
+A.3 Definition - How?
+A critical decision in this stage is whether to incorporate proposals in the annotation process.
+While proposals can significantly reduce annotation costs, they may introduce a bias (Jachimowicz et al., 2019; Schmarje et al., 2023).
+This bias results from annotators’ tendency to favor the proposed class, skewing the label distribution $P(L^{x}=\cdot)$ towards it.
+For making this decision, several parameters, such as the class confusion matrix $c$ and the expected annotation speedup $S$, need estimation or a defined approach for later approximation (see subsection B.2 and subsection B.3 for techniques).
+The first question to address is whether introducing a bias is acceptable.
+In cases where bias is tolerable, proposals should be used.
+If not, the expected speedup $S$ becomes the determining factor.
+Proposals are more advantageous with a higher speedup, as a higher speedup allows for easier reversal of the introduced bias with more annotations.
+(Schmarje et al., 2023) identifies a trade-off point for speedup around 3; below this, it’s advisable to avoid proposals due to the difficulty in reversing the bias within a reasonable budget.
+The class confusion matrix $c$ is utilized in post-processing (see subsection A.5) and not needed until then.
+A.4 Annotation Process
+This step focuses on creating the annotations, with various preparations depending on whether proposals are used.
+Literature indicates that proposals, particularly DC3 proposals, enhance annotation consistency, especially with ambiguous data (Papadopoulos et al., 2021; Desmond et al., 2021; Schmarje et al., 2021a).
+We recommend DC3 proposals (Schmarje et al., 2022b), which perform well with real-world ambiguous data and are compatible with many semi-supervised methods (Sohn et al., 2020; Li et al., 2020; Tarvainen and Valpola, 2017).
+Normal network predictions can also be used but might reduce overall performance.
+The labeled dataset $X_{l}$ serves as training data for a neural network to generate these proposals, which should be trained until reaching the same quality threshold $\mu$ as the annotators.
+Two important variables to define are the number of annotations required for consensus ($A_{\text{cons}}$) and the total number of annotations ($A^{\text{cons}}$).
+These variables balance between the best approximation of the ground-truth distribution $P(\hat{L}^{x}=\cdot)$ and the effort and cost of annotations.
+In subsection B.2 and subsection B.3, we discuss plausible numbers of annotations per hour and potential confidence guarantees for different combinations of $A_{\text{cons}}$ and $A^{\text{cons}}$.
+The annotations can either be created by different individuals or by showing the same images multiple times to one or multiple annotators.
+For the latter case, it is important to think about a cooldown period, the time between showing the same image again.
+This period can be quite small, down to a few hours, as long as there is a high probability that the annotator will not remember a case between annotations.
+A.5 Post-Processing
+The use of proposals in annotation can introduce biases due to the default effect (Jachimowicz et al., 2019).
 Schmarje et al.
-propose in [44] with CleverLabel a combination of classblending and bias correction to improve the approximated distribution, which should be applied if proposals were used during the annotation process.
-If no proposals were used to create the annotations, it is still recommended to improve the distribution $P(L^{x}=\cdot)$ by blending with the class distribution $c$.
-3 Evaluation
-The goal of this paper is to solve real-world application annotation tasks such as the medical task of vertrebral fracture classification.
-For this purpose, we will describe how the previously defined strategy can be applied to this task on the publicly available datasets Verse2019[35] and Verse2020[49].
-An analysis of the benefits or shortfalls of the decisions is given in subsection 3.2.
-3.1 Applying the Strategy
-We use the Verse datasets for better reproducibility of our osteoporotic vertrebral fracture classification.
-As in [35], we consider only at thoracolumbar vertebrae because osteoporotic fractures occur mainly in this region of the spine, as shown in Figure 2.
-For the classification of the fractures, we use an adaptation of the most commonly used score, the semi-quantitative Genant score [11].
-As shown in Figure 2, the score defines four classes, which are mainly guided by the height reduction of the vertebrae compared to its neighbors or unfractured vertebrae.
-However, degenerative deformities (e.g.
-short vertebral height) of up to 20% are ignored for the osteoporotic fracture classification.
-Partially, these degenerative cases can not be assessed in the dataset because the resolution is lower than in normal radiographs.
-As shown in Table 1, this results in 3,761 individual vertebrae with a high class imbalance towards class zero.
-Overall, we have at least 1.1% samples per class and can apply common semi- or supervised techniques.
-Figure 2: Illustration of a spine and definition of height reduction classes
-[…]
-[…]
-[…]
-[…]
+(2023) propose CleverLabel, a combination of class blending and bias correction, to improve the approximated distribution if proposals were used.
+If no proposals were employed, enhancing the distribution $P(L^{x}=\cdot)$ by blending with the class distribution $c$ is still recommended.
+Appendix B Implementation and approximation details
+This subsection provides more in-depth descriptions for some of the previous parts.
+They are separated to allow a concentration on the most important information without diluting it with necessary but minor considerations.
+We start this section with a description of additional minor consideration which apply to the complete annotation strategy (see subsection B.1).
+Additionally, we will discuss how the expected workload for the human annotators can be calculated (see subsection B.2) and what confidence with regard to the soft label can be expected given a certain number of annotations (see subsection B.3).
+Lastly, we will reiterate the most important information about CleverLabel from (Schmarje et al., 2023).
+B.1 Additional minor considerations for the Annotation Strategy
+  • All decisions should typically be supported by domain experts or the end users which will use the labeled data in the end.
+  • Constraints like collaboration modes and support for proposals should be considered.
+For example, if crowd-sourcing should be used for annotation, this restricts the appropriate platforms.
+Due to the fact the decision if proposals will be used is made later in subsection A.3, it is advisable to ensure that these future options are supported to avoid unnecessary backtracking.
+  • If the training of the annotators is not successful it might be necessary to step back and rethink previous decision to allow a successful training.
+For example, special training with experts could be held, more detailed description of the task could be given and general feedback of annotators could be considered.
+  • Proposals $\rho_{x}$ with a low ground truth probability ($P(\hat{L}^{x}=\rho_{x})=0$) should be avoided and (Schmarje et al., 2022b) indicates only up to 10% of all proposals should be of such poor quality to allow successful training.
+The distribution $P(\hat{L}^{x}=\cdot)$ might not be known at this point but if available this should be checked in addition to the threshold $\mu$.
+  • The variables $c,\delta$ and $S$ may be approximated in various ways.
+One possible approach is defined in (Schmarje et al., 2023) by annotating about 100 images at least 10 times for the estimation of $c$, $\delta$ and $P(\hat{L}^{x}=\cdot)$.
+The speedup $S$ can for example be measured on a small subset with and without proposals during the training of the annotators.
+However, if the variables are not required during the decision process described above, they could also be approximated on the real annotation data or ignored entirely.
+For example, $c$ can easily be approximated on the complete dataset if no proposals are used during the annotation process and an approximation of $S$ is not required if a small introduction of a bias is acceptable.
+The variables could also be approximated with an educated guess either based on previous knowledge or experiments.
+  • The question is how $A_{\text{cons}}$ and $A^{\text{cons}}$ should be selected.
+We concentrate on two major aspects, what is a feasible workload for the annotators and what guarantees can be given.
+See for more details the approximations in subsection B.2 and subsection B.3.
+  • The class blending (CB or CleverLabel) in the post-processing step may not be necessary if the number of annotations is high (e.g.
+above 50), since the distribution is already approximated quite accurately.
+B.2 Approximation of workload
+The expected workload can be approximated as the expected number of annotations per hour done by one annotator as $\frac{a}{h}$.
+Depending on the expected portion of complete consensus of the dataset $p_{c}$ and the total dataset size $N=|X_{u}|$, we can calculate the expected annotation time in hours with $A_{N}=N\cdot(p_{c}\cdot A_{\text{cons}}+(1-p_{c})\cdot A^{\text{cons}})/\frac{a}{h}$.
+In Figure 5 the required annotation time $A_{N}$ in days for $N=10,000,A_{\text{cons}}=10,A^{\text{cons}}=50$ with varying $p_{c}$ is given as well as values reported in (Schmarje et al., 2022b).
+The reported values for $\frac{a}{h}$ are between 1,000 and 4,500 without proposals or 3,000 and 10,000 with proposals.
+Thus, even in the worst-case the expected workdays with proposals are less than 10 days, or two weeks, which should be feasible in most cases especially when multiple annotators are available.
+These values are examples and have of course to be adopted to the specific use-case.
+Figure 5: Annotation time in days based on expected consensus ratio $p_{c}$ and annotations per hour $\frac{a}{h}$, reported values from (Schmarje et al., 2022b) and with 95% agreeing votes threshold $\hat{p_{c}}$
+B.3 Approximation of confidences
+The confidences of the calculated label distributions should be considered when $A^{\text{cons}}$ and $A_{\text{cons}}$ are defined.
+Considering the binary case of the annotation task for class $k$ for each image $x$, i.e.
+whether class $k$ is annotated for $x$ or not, and having a sufficient number of annotations $A$, the 95%-confidence interval can be calculated using the Wald confidence interval (Lyman, 1993).
+The interval is given by
+$\begin{split}P(\hat{L}^{x}=k)&\pm Z_{95\%}\cdot\sqrt{\frac{p_{k}(1-p_{k})}{A}}\end{split}$ (1)
+with 95%-confidence value, which is 1.96, and $p_{k}$ is the unknown true GT probability for class $k$.
+The confidence interval is the largest for $p=0.5$.
+The formula can be rearranged to calculate the required annotations $A$ if a 95% confidence interval around the approximation $P(\hat{L}^{x}=k)$ of width $W$ and margin $W/2$ is to be guaranteed.
+$\begin{split}\frac{W}{2}&=Z_{95\%}\cdot\sqrt{\frac{p_{k}(1-p_{k})}{A}}\\ \frac{W^{2}}{4}&=Z_{95\%}^{2}\cdot\frac{p_{k}(1-p_{k})}{A}\\ A&=\frac{4\cdot Z_{95\%}^{2}\cdot p_{k}(1-p_{k})}{W^{2}}\\ \end{split}$ (2)
+Figure 6 visualizes the required annotations $A$ for various $p$ and $W$.
+These calculations can guarantee with 95% confidence that the confidence interval has a width of $W\approx 1$ for $A=3$, $W\approx 0.62$ for $A=10$ and $W\approx 0.28$ for $A=50$.
+However, a high number of $A$ is indirectly assumed because the Wald confidence interval approximates the original binomial distribution with a normal distribution.
+This approximation is especially inaccurate for small numbers of $A$ and $p$ near 0 or 1, as pointed out by Lyman (1993).
+They suggest to approximate the 95% confidence intervals with $((0.25)^{\frac{1}{A}},1)$ when $p$ is close to one.
+This would give a lower bound of the interval of $0.63$ for $A=3$, $0.87$ for $A=10$ and $0.97$ for $A=50$.
+All these calculations assume an idealized version of the annotators, which must be taken into account in their interpretation.
+For this reason, it will not be investigated further how the confidence intervals can be approximated more precisely, but conclude that a consensus decision is uncertain for $A=3$ annotations, but reasonable for ten or more annotations.
+Similarly, approximations of $p$ with a confidence interval width less than 0.3 are only reasonable for about 50 annotations.
+Figure 6: Number of images for confidence width $W$, logarithmic scale
+B.4 Details on using CleverLabel and CB
+This section provides an overview about the methodology proposed in (Schmarje et al., 2023).
+These elements are reiterated here to keep this paper self-contained.
+Schmarje et al.
+presented in (Schmarje et al., 2023) two different approaches to improve the resulting distribution when using proposals.
+The approaches were called class blending (CB) and bias correction (BC) and formed together the method CleverLabel.
+Class blending uses a constant transition matrix $c$ between classes and blends this fixed class distribution with the calculated distribution.
+The constant transition matrix is motivated by the fact that for example a car and truck are more easily confused than a car and a dog.
+The bias correction used the theoretical analysis of the introduced bias when using proposals to invert this process.
+A main variable for this calculation is the dataset specific offset $\delta$.
+This offset describes how likely it is that a human will accept the proposal even if the proposed class would normally not been considered.
+A offset of zero means that the proposal has no influence while a value of one means that the proposal is always accepted regardless of the content.
+For details, please see the original publication (Schmarje et al., 2023).
+Appendix C Differences to previous work
+(a) CT slice
+(b) CT projection
+(c) Mask projection
+(d) Blended
 Figure 7: Illustration of different image modalities
-Unlike to the original author of the Verse2019 [35] dataset, we will not use medical experts for the annotation and will use 2D projections of the original 3D CT data.
+Unlike to the original author of the Verse2019 (Löffler et al., 2020) dataset, we will not use medical experts for the annotation and will use 2D projections of the original 3D CT data.
 Our decision to use non-medical experts has two main reasons.
@@
 Second, we plan to generate tens of thousands of annotations, which would not be feasible with medical experts, who are in high demand.
-While we will not use medical experts for the majority of the annotations, we have worked closely with the radiologist at XXXX (censored for review).
+While we will not use medical experts for the majority of the annotations, we have worked closely with the radiologist at University Medical Center Schleswig-Holstein.
 We discussed our design choices with them, particularly regarding the definition of the task and the representation of the vertebrae used.
 As mentioned above, we trained the hired workers on this specific task and ensured that they achieved a macro accuracy and F1 score of at least $\mu=0.6$ on the complete dataset with and without proposals for at least two scores before they could annotate the final scores.
-We do not use the original 3D CT data (see Figure 3) for the annotation but rather a 2D projection of the central slices for each vertebrae (see Figure 3) and its segmentation (see Figure 3).
+We do not use the original 3D CT data (see 7(a)) for the annotation but rather a 2D projection of the central slices for each vertebra (see 7(b)) and its segmentation (see 7(c)).
 Additionally, we show the vertebrae above and below the target vertebrae to allow an evaluation of the height reduction.
-A blended example is shown in Figure 3.
+A blended example is shown in 7(d).
 There are three major benefits/reasons of using 2D projections over 3D CT data.
 First, the Genant scores were defined on 2D X-Ray and 2D projections of CT data are visually very similar to X-Ray and thus are better comparable.
-Second, we want to use the reported methods from the literature for the proposal generation [46], which were only tested on 2D data.
-While it should be theoretically possible to extend the used networks to 3D, such major structural changes could lead to an inferior result which would dilute the results, such experiments may be future work.
+Second, we want to use the reported methods from the literature for the proposal generation (Schmarje et al., 2022b), which were only tested on 2D data.
+While it should be theoretically possible to extend the used networks to 3D, such major structural changes could lead to an inferior result which would dilute the results.
+But such experiments may be future work.
 Lastly, when it comes to diagnostics, radiologists mainly look at one central slice which we include and the major 3D structures can be detected on this projection.
+Annotators and annotation platform
 We hired 5 workers to annotate vertebrae in a web-based environment developed for this evaluation.
 Public annotation platforms do not support the combination of clusters and classes that we want to evaluate in this paper.
-Details about the platform are given in the supplementary.
-As $X_{l}$ we used the partially available labels from [35].
+Figure 8 displays the used annotation website.
+It is important to note, that all images are displayed to the user in a grid.
+In the case without proposal (see 8(a)), the shown classes are mixed and all have to be annotated manually by the user.
+In the second case with proposal (see 8(b)), most of the images are of the same class and can be annotated simultaneously, while only the errors need to be corrected.
+As $X_{l}$ we used the partially available labels from (Löffler et al., 2020).
+The software is a custom implementation to handle various inputs like classification and clusters. We plan to release a version in the future and please contact us if you are interested in a preview.
+(a) No Proposal
+(b) With DC3 proposal cluster
+Figure 8: Screenshots of the annotation website on the MiceBone dataset from (Schmarje et al., 2022a), The sidebar for data selection is hidden for data protection reasons.
+Appendix D Used parameters and their calculation
+D.1 Approximation of desired variables
 In a medical consensus process in an expert panel, the proposals of different people are unified, but influences between people due to social dynamics cannot be excluded.
@@
 To allow a validation of our decisions, we will also raise annotations without proposals and thus we can approximate the class distributions $c$ and speedup $S$ based on these annotations.
-Additional approaches to approximate $S$ as well as the approximation of a data-specific offset $\delta$ (see [44]) are discussed in the supplement.
+Additional approaches to approximate $S$ as well as the approximation of a data-specific offset $\delta$ (see (Schmarje et al., 2023)) are discussed in the supplement.
 Based on the training annotations, we estimated the $\frac{a}{h}$ to be about 2,000 and 3,000 without and with proposals, respectively.
-These results are similar values previously reported in [46].
-For $A_{c}=10$, $A^{c}=50$ and estimated $p_{c}=0.5$, this would result in about 250,000 annotations with and without proposals.
+These results are similar values previously reported in (Schmarje et al., 2022b).
+For $A_{\text{cons}}=10$, $A^{\text{cons}}=50$ and estimated $p_{c}=0.5$, this would result in about 250,000 annotations with and without proposals.
 Furthermore, based on the estimated $\frac{a}{h}$ of about 2,500, this resulted in a workload of about 20 hours per annotator which was within our given budget.
-So we chose the number of annotations at the recommended values of ${A_{c}=10,~{}A^{c}=50}$.
+So we chose the number of annotations at the recommended values of ${A_{\text{cons}}=10,\leavevmode\nobreak\ A^{\text{cons}}=50}$.
 Additionally, we planned to annotate the 3D data to evaluate the quality difference between 2D and 3D.
-We limited these annotations to the test set of the original Verse2019 dataset [35].
-3.2 Analysis
-The annotators completed more than 100 iterations of annotation, half of them with proposals and half without.
-However, one annotator did not meet the acceptance threshold $\mu$ and was excluded along with the training iterations, leaving 80 valid iterations or a total of about 250,000 annotations.
-Additionally, 11 iterations were performed using 3D data from the test dataset for comparison.
-The average annotation rate was calculated to be 3,259.36 $\frac{a}{h}$, which means that it takes about one hour to complete the 3,761 annotations with the additional technical overhead and pauses.
-The comparison of the resulting dataset-specific variables with previously reported values [46] is given in Table 1.
-We see that the dataset offset $\delta$ and the rate of near-consensus rate $\hat{p_{c}}$ are similar to previously reported results.
-It is important to note that $p_{c}=0.27$, which is significantly smaller, resulting in more data with $A^{c}$ number of annotations.
-It should be evaluated in the future, if instead of total consensus a high number of agreeing votes is sufficient to reduce the cost of the annotation process.
-The largest class is greater than in the previous reports and the speedup $S$ is smaller in Table 1.
-This relationship is theoretically justified because a proposal is less important since the majority of images already belongs to a class and thus defines some kind of a proposal.
-name $\delta$ $\hat{p_{c}}$ classes largest class [%] $S$
-Verse (ours) 0.1143 0.6833 4 90.11 1.1636
-CIFA10-H 0.0973 0.7766 10 10.16 2.4665
-MiceBone 0.3636 0.4775 3 70.48 1.4471
-Plankton 0.5784 0.7368 10 30.37 4.4319
-Turkey 0.2164 0.6863 3 75.95 1.4877
-Table 1: Comparison of dataset-specific variables with previous reported values [44]$\delta$ is the data specific offset when annotating with proposals described in [44].
-$S$ is the speedup between annotating with and without proposals.
-$\hat{p_{c}}$ describes the percentage of data with at least 95% consistent annotations.
+We limited these annotations to the test set of the original Verse2019 dataset (Löffler et al., 2020).
+We used annotations created during the training of the annotators to estimate $\delta$.
+We randomly picked 20 examples without consensus from the annotations without proposals.
+Based on these samples, the proposal per image $p_{x}$, the probabilities $P(L_{x}=\rho_{x})$ with and without proposal and the formula for simulated proposal acceptance of (Schmarje et al., 2023), we calculated $\delta\approx 0.11$.
+D.2 Used hyperparameters
+The used hyperparameters are defined in Table 2. For more detail see the respective source code 111 https://github.com/Emprime/dc3, https://github.com/Emprime/dcic .
+Table 2: Used hyperparameters for the proposal and evaluation network, $\dagger$ approximated based on number of image iterations, epochs not counted
+Parameter Proposal network Evaluation network
+Architecture DC3(Schmarje et al., 2022b) + Mean-Teacher(Tarvainen and Valpola, 2017) ResNet50(He et al., 2016)
+Learning rate 0.03 0.1
+Epochs 600$\dagger$ 60
+Optimizer SGD SGD
+Batch size 64 128
+Weight Decay 0.0005 0.0005
+Pretrained Weights N/A ImageNet
+Source code (Schmarje et al., 2022b) (Schmarje et al., 2022a)
+Appendix E Additional results of analysis
+E.1 Learning effect
+Regarding the learning effect of the annotators, the training iterations showed an increase in F1 scores over time.
+The improvement in F1 scores across all annotators between the training and the last three iterations was 6.79 without proposals and 9.19 with proposals.
+When comparing the benefit of DC3 proposals versus no proposals, the average improvement in F1 scores was 3.27 for the first three iterations and 0.41 for the last three iterations.
+The proposal still has a benefit in the end but the annotators can more successfully annotate event without them.
+Moreover, the introduction of proposal-based annotations resulted in a reduction in training time which also diminished over time.
+The first three iterations saw a reduction of 13.49 minutes, while the last three iterations experienced a decrease of 4.55 minutes.
+Overall, we can conclude that the proposals helped to improve the quality and reduce the annotation time.
+However, this effect was diminished over time while the performance with and without proposals increases.
+Thus, we see a learning effect of the annotators.
+E.2 Evaluation of human performance
+Table 3: Comparison of human and network performance on Verse2019 (Löffler et al., 2020) – * based on ten randomly sampled annotations
 $F1$
@@
 Proposal Network 0.57
-Wei et. al [59] 0.59
+Wei et al. (Wei et al., 2022) 0.59
 Ours* 0.63
-Table 2: Comparison human and network performance on Verse2019 [35] – * based on ten random sampled annotations
-A comparison of the annotator and network performance on the Verse2019 [35] test set is given in Table 2.
-When comparing the macro F1 scores in 2D between annotations performed with and without proposals and those with proposal-based guidance achieved a significantly higher F1 score of about 2%.
-This means that proposal-based guidance led to better annotations by the annotators.
-The average macro F1 score for 3D annotations shows no significant difference compared to the 2D annotations, meaning that the modality difference was of no to little importance to the annotators.
+A comparison of the annotator and network performance on the Verse2019 (Löffler et al., 2020) test set is given in Table 3.
+When comparing the macro F1 scores in 2D between annotations performed with and without proposals and those with proposal-based guidance, the latter achieved a significantly higher F1 score of about 2%.
+This means that the employment of proposal-based guidance resulted in improved annotations by the annotators.
+The average macro F1 score for 3D annotations shows no significant difference compared to the 2D annotations, meaning that the modality difference was of no to little importance to the annotators as expected in subsection 3.1.
 The marginal class distribution on all vertebrae from the test set between the ground truth and the majority vote for 2D annotations was quite similar, with percentages of 80.33, 10.38, 6.01, and 3.28 versus 80.33, 10.93, 5.74, and 3.01, respectively.
@@
 These results indicate that the annotators produced annotations of similar quality to the original test data, with higher uncertainties for potentially difficult or incorrect majority votes.
-The best performing network on the validation data for the proposal generation did not surpass the set target threshold $\mu=0.6$ with a macro F1 of 0.57.
-However, it surpassed $\mu$ on the validation and training data and it was decided that the performance was close enough.
-The proposal network was a semi-supervised Mean-Teacher model [53] with overclustering and default hyperparameters from [46].
-The performance of the model was significantly better when trained only on the mask (see Figure 3) with only the desired vertebra visible instead of all other 2D modalities.
-We attribute this difference to the fact that the model could be more easily confused with vertebrae that are not currently of interest and to the more complicated representation of information e.g.
-in the form of blended images.
-In terms of the performance of the network trained on our new annotated data, but evaluated on the original test ground truth, we found that more annotations sampled from the annotated data generally led to better performance.
-This sampling from the human annotation could achieve random peaks up to about 0.7 F1-score.
-A standard ResNet50 model [18] with default hyperparameters (see supplementary for details) and cross-entropy as loss function could achieve a score of 0.63, which is higher than the more complex proposal network (trained on the original training data) and previously reported values by Wei et al.
-for non-specialized networks [59].
-While they show that they can achieve higher scores with more problem-specific loss functions, with our data-centric approach, we can achieve an improvement without changing the model or training purely from the input data.
-Future potential improvement to also use special loss functions and better backend-models non-withstanding.
-method number of annotations
-proposal blending correction 1 3 5 10
-DC3 balanced yes 0.4481 $\pm$ 0.1957 (0.00%) 0.2425 $\pm$ 0.0741 (0.00%) 0.1832 $\pm$ 0.0095 (0.00%) 0.1615 $\pm$ 0.0129 (0.00%)
-DC3 balanced no 1.0309 $\pm$ 0.1750 (130.07%) 0.2704 $\pm$ 0.0760 (11.53%) 0.2105 $\pm$ 0.0256 (14.85%) 0.4546 $\pm$ 0.4343 (181.49%)
-DC3 no yes 1.7382 $\pm$ 0.1553 (287.91%) 0.2499 $\pm$ 0.0374 (3.08%) 0.2081 $\pm$ 0.0171 (13.57%) 0.2798 $\pm$ 0.1325 (73.25%)
-no only blends no 0.2311 $\pm$ 0.0224 (-48.42%) 0.2394 $\pm$ 0.0930 (-1.27%) 0.1904 $\pm$ 0.0209 (3.88%) 0.1644 $\pm$ 0.0062 (1.79%)
-no balanced no 0.8565 $\pm$ 0.4466 (91.14%) 0.1966 $\pm$ 0.0245 (-18.90%) 0.1678 $\pm$ 0.0224 (-8.44%) 0.1568 $\pm$ 0.0165 (-2.92%)
-no no no 0.5578 $\pm$ 0.1259 (24.49%) 0.2451 $\pm$ 0.0127 (1.10%) 0.5435 $\pm$ 0.4285 (196.61%) 0.1898 $\pm$ 0.0380 (17.52%)
-Table 3: Resulting KL across different methods and number of annotations – The first method is the recommendation based on our defined guidelines.
-The next two methods use proposals and the last three methods do not use proposals which means their annotation time is also slower in comparison to the recommendation.
-Thus, one should not compare with the same number of annotations but higher number of annotations for DC3 proposals because the higher number can be achieved potentially in the same or less time.
-The results are given as KL $\pm$ STD ( Relative Change in % with regard to recommended method)
-In Table 3, we compare the network predictions with the approximated soft ground truth distribution ($P(L^{x}=\cdot)$) using the Kullback-Leibler-divergence [27].
-Comparing the use of DC3 proposals with and without blending or correction, we see that the results are better with both improvement methods proposed in [44].
-Thus, we can confirm that the application of CleverLabel (the combination of both improvements) should be done in the post-processing step.
-We can also confirm that applying balanced blending to annotations without proposals improves the results.
-Using only the class blending based on the majority class seems to work particularly well for this task.
-However, as the number of annotations increases, the benefit decreases and from 5 annotations onward, it becomes worse than the recommended method of our defined strategy.
-In our description of how to apply the proposed annotation strategy (see subsection 3.1), we stated that the introduction of a bias is not a major concern.
-Nevertheless, we can validate that the strategy leads to the best results even if a bias would not be acceptable.
-Based on the speedup of about 1.2, the recommendation would have been to not annotate with proposals and to use only balanced blending, which is the best method for multiple annotations.
-Theoretically, if we had a high speedup, we could compare a higher number of annotations with proposals to fewer annotations without proposals.
-For example, a speedup at the decision boundary of 3, 5 annotations with proposal and 3 annotations without proposal, we have a lower annotation cost based on the calculations form [44] (5 $\cdot$ 0.33 vs.
-3) and have a lower KL score.
-Thus, even in the case of higher speedups, our strategy suggests the correct choice.
-4 Conclusion
-In conclusion, this study presents a general strategy for annotating and processing ambiguous images with verification in vertebral fracture diagnosis.
-Overall, the human performance analysis demonstrated the effectiveness of proposal-guided annotation, resulting in improved F1 scores and reduced training time.
-A neural network trained on newly annotated data achieved a higher classification performance on the original test set and can potentially more than halve the KL score on our new baseline.
-Based on the results for all possible scenarios, we were able to verify the effectiveness of our strategy based on the literature for the given use case.
-By providing practical guidance and verifying its application on a real-world vertebral fracture dataset, this research contributes to the creation of high-quality datasets and advances the field of data-centric deep learning.
-Supplementary Material
-Approximation of workload
-We look at the expected number of annotations per hour done by one annotator as $\frac{a}{h}$.
-Depending on the expected portion of complete consensus of the dataset $p_{c}$ and the total dataset size $N=|X_{u}|$, we can calculate the expected annotation time in hours with $A_{N}=N\cdot(p_{c}\cdot A_{c}+(1-p_{c})\cdot A^{c})/\frac{a}{h}$.
-In Figure 4 the required annotation time $A_{N}$ in days for $N=10,000,A_{c}=10,A^{c}=50$ with varying $p_{c}$ is given as well as values reported in [46].
-The reported values for $\frac{a}{h}$ are between 1,000 and 4,500 without proposals or 3,000 and 10,000 with proposals.
-Thus, even in the worst-case the expected workdays with proposals are less than 10 days, or two weeks, which should be feasible in most cases especially when multiple annotators are available.
-These values are examples and have of course to be adopted to the specific use-case.
-Figure 4: Annotation time in days based on expected consensus ratio $p_{c}$ and annotations per hour $\frac{a}{h}$, reported values from [46] and with 95% agreeing votes threshold $\hat{p_{c}}$
-Approximation of confidences
-Considering the binary case of the annotation task for class $k$ for each image $x$, i.e.
-whether class $k$ is annotated for $x$ or not, and having a sufficient number of annotations $A$, the 95%-confidence interval can be calculated using the Wald confidence interval [37].
-The interval is given by
-$\begin{split}P(L^{x}=k)&\pm Z_{95\%}\cdot\sqrt{\frac{p_{k}(1-p_{k})}{A}}\end{split}$ (1)
-with 95%-confidence value, which is 1.96, and $p_{k}$ is the unknown true GT probability for class $k$.
-The confidence interval is the largest for $p=0.5$.
-The formula can be rearranged to calculate the required annotations $A$ if a 95% confidence interval around the approximation $P(L^{x}=k)$ of width $W$ and margin $W/2$ is to be guaranteed.
-$\begin{split}\frac{W}{2}&=Z_{95\%}\cdot\sqrt{\frac{p_{k}(1-p_{k})}{A}}\\ \frac{W^{2}}{4}&=Z_{95\%}^{2}\cdot\frac{p_{k}(1-p_{k})}{A}\\ A&=\frac{4\cdot Z_{95\%}^{2}\cdot p_{k}(1-p_{k})}{W^{2}}\\ \end{split}$ (2)
-Figure 5 visualizes the required annotations $A$ for various $p$ and $W$.
-These calculations can guarantee with 95% confidence that the confidence interval has a width of $W\approx 1$ for $A=3$, $W\approx 0.62$ for $A=10$ and $W\approx 0.28$ for $A=50$.
-However, a high number of $A$ is indirectly assumed because the Wald confidence interval approximates the original binomial distribution with a normal distribution.
-This approximation is especially inaccurate for small numbers of $A$ and $p$ near 0 or 1, as pointed out by Lyman et al.
-[37].
-They suggest to approximate the 95% confidence intervals with $((0.25)^{\frac{1}{A}},1)$ when $p$ is close to one.
-This would give a lower bound of the interval of $0.63$ for $A=3$, $0.87$ for $A=10$ and $0.97$ for $A=50$.
-All these calculations assume an idealized version of the annotators, which must be taken into account in their interpretation.
-For this reason, it will not be investigated further how the confidence intervals can be approximated more precisely, but conclude that a consensus decision is uncertain for $A=3$ annotations, but reasonable for ten or more annotations.
-Similarly, approximations of $p$ with a confidence interval width less than 0.3 are only reasonable for about 50 annotations.
-Figure 5: Number of images for confidence width $W$, logarithmic scale
-Additional minor considerations for Annotation Strategy
-  • All decisions should typically be supported by domain experts or the end users which will use the labeled data in the end.
-  • Constraints like collaboration modes and support for proposals should be considered.
-For example, if crowd-sourcing should be used for annotation, this restricts the appropriate platforms.
-Due to the fact if proposals will be used is decided later in subsection 2.3, it is advisable to ensure that these options are supported to avoid unnecessary backtracking.
-  • If the training of the annotators is not successful it might be necessary to step back and rethink previous decision to allow a successful training.
-For example, special training with experts could be held, more detailed description of the task could be given and general feedback of annotators could be considered.
-  • Proposals $\rho_{x}$ with a low ground truth probability ($P(L^{x}=\rho_{x})=0$) should be avoided and [46] indicates only up to 10% of all proposals should be of such poor quality to allow successful training.
-The distribution $P(L^{x}=\cdot)$ might not be known at this point but if available this should be checked in addition to the threshold $\mu$.
-  • The variables $c,\delta$ and $S$ may be approximated on various ways.
-One possible approach is defined in [44] by annotating about 100 images at least 10 times for the estimation of $c$, $\delta$ and $P(L^{x}=\cdot)$.
-The speedup $S$ can for example be measured on a small subset with and without proposals during the training of the annotators.
-However, if the variables are not required during the decision process described above, they could also be approximated on the real annotation data or ignored entirely.
-For example, $c$, can easily be approximated on the complete dataset if no proposals are used during the annotation process and an approximation of $S$ is not required if a small introduction of a bias is acceptable.
-The variables could also be approximated with an educated guess either based on previous knowledge or experiments.
-  • The question is how $A_{c},A^{c}$ should be selected.
-We concentrate on two major aspects, what is a feasible workload for the annotators and what guarantees can be given.
-See for more details the approximations in Approximation of workload and Approximation of confidences.
-  • The blending step may not be necessary if the number of annotations is high (e.g. above 50), since the distribution is already approximated quite accurately.
-Approximation of $\delta$
-We used annotations created during the training of the annotators to estimate $\delta$.
-We randomly picked 20 examples without consensus from the annotations without proposals.
-Based on these samples, the proposal per image $p_{x}$, the probabilities $P(L_{x}=\rho_{x})$ with and without proposal and the formula for simulated proposal acceptance of [44], we calculated $\delta\approx 0.11$.
-Used hyperparameters
-The used hyperparameters are defined in Table 4. For more detail see the respective modified source code.
-Parameter Proposal network Evaluation network
-Architecture DC3[46] + Mean-Teacher[53] ResNet50[18]
-Learning rate 0.03 0.1
-Epochs 600$\dagger$ 60
-Optimizer SGD SGD
-Batch size 64 128
-Weight Decay 0.0005 0.0005
-Pretrained Weights N/A ImageNet
-Source code [46] [45]
-Table 4: Used hyperparameters for the proposal and evaluation network, $\dagger$ approximated based on number of image iterations, epochs not counted
-Additional results of the learning effect
-Regarding the learning effect of the annotators, the training iterations showed an increase in F1 scores over time.
-The improvement in F1 scores across all annotators between the training and the last three iterations was 6.79 without proposals and 9.19 with proposals.
-When comparing the benefit of DC3 proposals versus no proposals, the average improvement in F1 scores was 3.27 for the first three iterations and 0.41 for the last three iterations.
-The proposal still has in the end a benefit but the annotators can more successfully annotate event without them.
-Moreover, the introduction of proposal-based annotations resulted in a reduction in training time which also diminished over time.
-The first three iterations saw a reduction of 13.49 minutes, while the last three iterations experienced a decrease of 4.55 minutes.
-Overall, we can conclude that the proposals helped to improve the quality and reduce the annotation time.
-However, this effect was diminished overtime while the performance with and without proposals increase.
-Thus, we see a learning effect of the annotators.
-Annotation platform
-Figure 6 displays the used annotation website.
-It is important to note, that all images are displayed to the user in a grid.
-In the case without proposal (see Figure 6), the shown classes are mixed and all have to be annotated manually by the user.
-In the second case with proposal (see Figure 6), most of the images are of the same class and can be annotated simultaneously, while only the errors need to be corrected.
-[…]
-[…]
-Figure 6: Screenshots of the annotation website on the MiceBone dataset from [45], The sidebar for data selection is hidden for data protection reasons and to keep the anonymity.
+Appendix F Dataset Card
+Attribute Details
+Dataset Name VerSe 2D Projections with Soft Labels
+Dataset URL https://zenodo.org/records/8115942
+Original datasets:
+VerSe’19: https://osf.io/nqjyw/ VerSe’20: https://osf.io/t98fz/
+Description The dataset contains 2D projection of the central slices for vertebra of the VerSe (Large Scale Vertebrae Segmentation Challenge) datasets.
+These are collections of multi-detector CT scans for vertebral labelling and segmentation, significant for its size, diversity, and detailed annotations.
+The data, collected in conjunction with MICCAI 2019 and 2020, is anonymized and includes a variety of normal and pathological cases.
+Data Collection Domain: Medical Imaging, 2D Projections of Spine Multi-detector Computed Tomography (MDCT)
+Annotation Details: Soft-labels by evaluating proposal-based human non-expert annotations
+VerSe’19: 160 CT scans from 141 patients, 1725 annotated vertebrae
+VerSe’20: 374 CT scans from 335 patients, 4505 annotations
+Data Preprocessing Scans and segementation masks have been anonymized and converted to NIFTI format (https://nifti.nimh.nih.gov/) in the VerSe dataset.
+The resolution of the CT images was adjusted to limit computational demands for deep learning algorithms.
+Coordinates of vertebral body centroids per vertebral level are stored in JSON format.
+Annotations The proposed multi-step algorithmic framework has been applied to generate annotations with or without proposals by trained non-experts, including post-processing with CleverLabel and class blending.
+Usage Development and evaluation of algorithms for vertebral labelling and segmentation in spine CT scans.
+VerSe‘19 and VerSe‘20 were used in challenges to evaluate accuracy, precision, and the effectiveness of segmentation and fracture detection as part of MICCAI 2019 and 2020.
+Known Limitations While including multiple sites, the datasets might still not represent all possible variations in spine pathology and patient demographics.
+VerSe’19: Exclusion of patients younger than 30 years and certain spinal conditions like bone metastasis
+Ethical Considerations Both VerSe studies were approved by the local institutional review board, which conducted a thorough review of the research methodology, ensuring adherence to ethical guidelines.
+All patient data included in this study were anonymized to protect the privacy and confidentiality of the individuals.
+Acknowledgements VerSe has been supported by the European Research Council under Horizon2020 (GA637164–iBack–ERC–2014–STG) and NVIDIA.
+Table 4: Overview of the VerSe 2D Projections with Soft Labels